### v1

# Ergodic Theorems and Equivalence of Green’s Kernel for Random Walks in Random Environments

Abstract
We study the Ergodic Properties of Random Walks in stationary ergodic environments without uniform ellipticity under a minimal assumption. There are two main components in our work. The first step is to adopt the arguments of Lawler to first prove a uniqueness principle. We use a more general definition of environments using Environment Functions. As a corollary, we can deduce an invariance principle under these assumptions for balanced environments under some assumptions. We also use the uniqueness principle to show that any balanced, elliptic random walk must have the same transience behaviour as the simple symmetric random walk. The second is to transfer the results we deduce in balanced environments to general ergodic environments(under some assumptions) using a control technique to derive a measure under which the local process is stationary and ergodic. As a consequence of our results, we deduce the Law of Large Numbers for the Random Walk and an Invariance Principle under our assumptions.
1. Introduction
The purpose of the discussion is to derive Ergodic Theorems and related results in the context of Random Walk in Random Environments(RWRE). This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers. Roughly speaking, it says there is a deterministic limiting velocity of the walk, and the second question is about the probability of the walk escaping to infinity along a given direction. Both have been studied in detail in $d=1$ by Solomon[11] in 1975. In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
Informally speaking, the problem regarding the Law of Large Numbers in RWRE is the following:
Conjecture 1.1.
Does there exist a deterministic vector $v$ such that
$\frac{X_{n}}{n}\to v\ \ \ \ \text{a.e.}$
for almost every choice of environment?
Varadhan in [12] showed that the sequence $X_{n}/n$ has at most two deterministic limit points $v_{1},v_{2}$ and if $v_{1}\neq v_{2}$ then there is a constant $a\geq 0$ such that $v_{1}=-av_{2}$. Furthermore, Nicholas Berger in [1] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0. These results assume that the environment is uniformly elliptic and i.i.d.. In our present discussion, we answer 1.1 in the affirmative for stationary ergodic nearest neighbour RWRE and show the existence of the limiting velocity under some assumptions.
We shall work under a more general definition of environments using Environment Function, which we shall introduce below. We shall establish under some minimal assumption that this chain has a stationary measure for balanced environments, adapting arguments from [7] for any environment function. This shall also provide an easy proof of Theorem 1 of [4]. This shall be derived by first establishing a uniqueness principle for concave functions on the lattice satisfying certain properties. We answer the following natural question for balanced random walks, which can be thought of for now as a random walk having equal probabilities of going forward or backward along each coordinate axis.
Conjecture 1.2.
Does every balanced, elliptic(probability of going along every axis at every point is positive) walk in $\mathbb{Z}^{d}$ recurrent in $d=1,2$ and transient in $d\geq 3$?
We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest. Using the uniqueness we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks. It is important to note that this section shall not deal with the notion of a Random Environment. This section is one of the most fundamental contributions, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
Our other main contribution will be to deduce the ergodicity of a certain stochastic process, which we will call the local process. This process chain is a projection at the origin of the “Environment Viewed from the Particle” Markov Chain or, as we will call it in this paper, the environment process with a suitable environment function. We consider the local process as environment process does not always possess suitable ergodic properties. This can be noted for Random Walks in Dirichlet Environments as studied by Christopher Sabot in [10].
Our definition of environments allows us to transfer our results from the balanced case to the general environment, by coupling the random walk and a reflected walk. We then combine a novel technique which we shall name as the Action of Linear Transformations and another control technique to execute a proof.
As a consequence of our ergodic theorem, we shall also be able to deduce an invariance principle for general environments. To the best of our knowledge, these theorems are novel. Our results hold for all uniformly elliptic RWRE, and the conditions we impose study the extent to which the condition of uniform ellipticity can be dropped.
2. Prerequisites
Before stating the main results, we formally define the setting we work under. We adopt the formalism as in [7].
Define by $\mathbb{Z}^{d}$ the integer lattice and $\{e_{i}\}_{i=1,..,d}$ the unit vectors. Define the set
$V_{d}=\{\pm e_{i}:i=1,\dots,d\}\cup\{0\}$
Define the following sets:
$S(d)$ $=\{(p_{0},p_{1},\dots,p_{2d}):p_{i}\geq 0,\ p_{0}+\sum_{i=1}^{2d}p_{i}=1\}$
$\mathcal{S}(d)$ $=\{\text{set of functions from }\mathbb{Z}^{d}\to S(d)\}.$
Let $\mathcal{B}(d)$, $\mathcal{B}(d)^{\prime}$ be the Borel-$\sigma$ algebra on $S(d)$, $\mathcal{S}(d)$ respectively.
We call any measurable $\mathscr{E}:\mathcal{S}(d)\to\mathcal{S}(d)$ as the environment function. This function essentially contains the information on the transition probabilities of the random walk based on the function $\omega$. This is essentially a mild but powerful generalisation of the environment, as we shall see.
Definition 2.1 (Environment).
Given an environment function $\mathscr{E}$, $\omega\in\mathcal{S}(d)$, we call the pair $(\omega,\mathscr{E})$ as the environment.
Let $\mu$ be a probability measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. We call the environment function balanced if
$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega(x))=\mathscr{E}_{i+d}(\omega(x))\quad i=1,2,...,d.$
We shall call the environment elliptic if
$\mu[\mathscr{E}_{k}(\omega)(x)>0:\forall x\in\mathbb{Z}^{d},\forall k]=1$
Define the group of shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$ on $\omega\in\mathcal{S}(d)$ given by
$(\tau_{x}\omega)(y)=\omega(x+y),\quad\forall x,y\in\mathbb{Z}^{d}.$
One notation we shall use interchangeably is the following:
Notation 1.
For $\mathscr{E}(\omega)\in\mathcal{S}(d)$, $v\in V_{d}$, $\mathscr{E}$ an environment function
$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}_{i}(\omega)(x)\quad\text{if}\penalty 10000\ v=e_{i}$
$\mathscr{E}(\omega)(x,-v)$ $=\mathscr{E}_{i+d}(\omega)(x)\quad\text{if}\penalty 10000\ v=-e_{i}$
$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}_{0}(\omega)\quad\text{if}\penalty 10000\ v=0.$
We shall assume $\mu$ to be stationary and ergodic with respect to the group of shifts $\tau_{x}$. More precisely,
$\mu[\tau_{-x}\omega\in A]=\mu[\omega\in A]\quad\forall A\in\mathcal{B}(d)^{\prime},\forall x\in\mathbb{Z}^{d}$
As a special case, $\mu$ is said to be an independent and identically distributed (i.i.d.) law if $\{\omega(x)\}_{x\in\mathbb{Z}^{d}}$ is an i.i.d. sequence of random variables under the measure $\mu$. One could think of this as independently assigning a point from $S(d)$ at every site.
Fix $\omega\in\mathcal{S}(d)$ and an environment function $\mathscr{E}$. Define the Markov Chain of the Random Walk $\{X_{\mathscr{E}(\omega)}(k)\}_{k\geq 0}$ in $\mathbb{Z}^{d}$ by the transition probabilities,
$\mathbb{P}[X_{\mathscr{E}(\omega)}(k+1)$ $=x+v|X_{\mathscr{E}(\omega)}(k)=x]=\mathscr{E}(x,v)\quad\forall v\in V_{d}$
Therefore, the generators of this Markov Chain is given by
$L_{\mathscr{E}(\omega)}(g(x))=\sum_{v\in V_{d}}\mathscr{E}(\omega)(x,v)(g(x+v)-g(x)),$
respectively for every $g:\mathbb{Z}^{d}\to\mathbb{R}$.
Define the environment process as the Markov Chain given by $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with the transition kernel given by
$K_{d}^{\mathscr{E}}(\omega,B)=\sum_{v\in V_{d}}{\mathscr{E}(\omega)}(o,v)\delta_{\tau_{v}\omega}(B)\quad\forall B\in\mathcal{B}(d)^{\prime}.$
We also define a projection operator $\pi_{E}$ for $E\subset\mathbb{Z}^{d}$,
$\pi_{E}(\omega)=\{\omega(x):x\in E\}.$
Then we define the local process associated with any discrete time Markov Chain $\{Y_{k}\}_{k\geq 0}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ as $\{\pi_{0}(Y_{k})\}_{k\geq 0}$.
Denote by $\mathbb{P}^{\mathscr{E}({\omega})}_{x}$: the law on path of the random walk started at $x$ in the environment $(\omega,\mathscr{E})$. Precisely, $\mathbb{P}^{{\mathscr{E}(\omega)}}_{x}$ is a measure on $((\mathbb{Z}^{d})^{\mathbb{N}},\mathcal{G)}$ where $\mathcal{G}$ is the $\sigma$-algebra generated by cylinder sets.
Therefore if ${\mathscr{E}(\omega)}\in\mathcal{S}(d)$ one has that
$\mathbb{P}^{{\mathscr{E}(\omega)}}_{x}(G):\mathcal{S}(d)\to[0,1]$
is measurable $\forall G\in\mathcal{G}$. Hence, we can define the annealed law to be the probability measure on $(\mathcal{S}(d)\times(\mathbb{Z}^{d})^{\mathbb{N}},\mathcal{B}(d)^{\prime}\times\mathcal{G)}$ as
$\mathbb{P}^{ann}_{x}(F\times G)=\int_{F}\mathbb{P}_{x}^{{\mathscr{E}(\omega)}}(G)d\mu(\omega)\quad\forall F\in\mathcal{B}(d)^{\prime},G\in\mathcal{G}.$
Denote the annealed law as $\mu^{x}=\mu\rtimes{\mathbb{P}}_{x}^{{\mathscr{E}(\omega)}}$.
Also, we shall define the drift of an environment at $x\in\mathbb{Z}^{d}$ as
$d_{\mathscr{E}}(x,\omega)=E[X_{{\mathscr{E}(\omega)}}(1)-x|X_{{\mathscr{E}(\omega)}}(0)=x].$
We make the following assumptions on the environment that is
Assumption 2.2.
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d))$. Define the quantity $c(\mathscr{E},\omega)(x)$ as follows
$c(\mathscr{E},\omega)(x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}$
Then we assume that
$\int_{\mathcal{S}(d)}c^{-p}d\mu<\infty.$
for some $p>d$.
Assumption 2.3.
Environment, or the balanced environment, is elliptic.
In our proofs we use the notation $|x|_{1}$ to denote the $L_{1}$ norm with respect to the counting measure in $\mathbb{R}^{k}$ given by
$|x|_{1}=\sum_{j=1}^{k}|x|_{j}.$
We use the notion of concave functions defined on the compact $L_{1}$ ball $B_{1}[0,r]\cap\mathbb{Z}^{d}$. A function $z:B_{1}[0,r]\cap\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ $(r>0)$ is said to be concave if
$z(x+v)+z(x-v)-2z(x)<0\quad x\in\operatorname{int}B_{1}[0,r]\cap\mathbb{Z}^{d}\quad\forall v\in V_{d}\setminus\{0\}$
where $\operatorname{int}$ denotes the topological interior of a set.
Also, note that we use the notion of mutual absolute continuity of two measures.
Definition 2.4.
Two measures $\mu,\nu$ on a measure space $(\Omega,\mathcal{F})$ are said to be mutually absolutely continuous if
$\forall A\in\mathcal{F}\quad\mu(A)=0\iff\nu(A)=0.$
We shall also use the notion of the quotient group, that is, given a group $G$ and an equivalence relation $R$ on it, we define the quotient group G’ as
$G^{\prime}=G/R.$
This is the group of all equivalence classes under the relation $R$. We shall denote the equivalence class of $g\in G$ by $g]_{R}$.
3. Main Results and Discussions
We first state the result that allows us to effectively deal with environments beyond uniform ellipticity. We shall demonstrate this through another result later in this section.
Theorem 3.1.
Let $d\geq 2$. Assume $\mu$ is a stationary, ergodic Law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a balanced environment function satisfying assumptions 2.2 and 2.3.
Then there exists a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ such that the environment process $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ is stationary, ergodic under this measure and is mutually absolutely continuous to $\mu$.
An important corollary of this result, which we shall also state but is not directly related to our discussion, is
Corollary 3.2.
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ and $\mathscr{E}$ a measurable balanced environment function. Fix $\omega\in\mathcal{S}(d)$. Then under assumptions 2.2 and 2.3,
$\frac{X_{\mathscr{E}(\omega)}([nt])}{\sqrt{n}}$
converges weakly to the standard Brownian Motion on $\mathbb{R}^{d}$ with non degenerate covariance matrix of the form $(b_{i}\delta_{ij})_{1\leq i,j\leq d}$ for $\omega$-$\mu$ a.e.
These results notably weaken the assumptions of ellipticity and derives an alternate proof of Theorem 1 of [5] for $d\geq 2$, which states that
Theorem 3.3 (Theorem 1(i) of [5]).
Let
$c(x)=\prod_{i=1}^{d}[\omega(x,e_{i})]^{\frac{1}{d}}.$
If the stationary, ergodic, and elliptic environment( $\omega(x,e_{i})=\omega(x,-e_{i})$ with $\mathscr{E}=\operatorname{Id}$) satisfies the condition
$\mathbb{E}[c(o)^{-p}]<\infty$
for some $p>d\geq 2$, then Corollary 3.2 holds.
To prove the above results, we derive a uniqueness theorem for concave functions on $\mathbb{Z}^{d}$, which also gives the following theorem on the transcience of balanced elliptic walks. The result compares the walk to the simple symmetric random walk on $\mathbb{Z}^{d}$ given by the law
$\mathbb{P}[X_{1}=x+v\Big|X_{0}=x]=\frac{1}{2d}.$
To state our result, we recall the so-called Green’s function for a random walk on $\mathbb{Z}^{d}$.
Definition 3.4.
Let $(\omega,\mathscr{E})$ be a balanced environment in $\mathbb{Z}^{d}$($d\geq 1$). Then the Green’s function $G^{\mathscr{E}(\omega)}:\mathbb{Z}^{d}\times\mathbb{Z}^{d}\to\mathbb{R}^{+}\cup\{0\}$ is given by
$\mathbb{E}_{x}[\sum_{j=0}^{\infty}\boldsymbol{1}_{X_{\mathscr{E}(\omega)}=y}]$
where the expectation is taken conditional on $X_{\mathscr{E}(\omega)}(0)=x$. For the simple symmetric random walk, we shall denote its Green’s function as just $G$.
Theorem 3.5.
Let $d\geq 1$. Let $(\omega,\mathscr{E})$ be any balanced elliptic environment. Then
$\frac{1}{2d}G(x,x)\leq G^{\mathscr{E}(\omega)}(x,x)\leq\frac{1}{c(\mathscr{E},x)^{d}}G(x,x)$
This shows the conclusions of Pölya’s Theorem, which states
Theorem 3.6 (Pölya).
The simple symmetric Random Walk on $\mathbb{Z}^{d}$ is transient in $d\geq 3$ and recurrent in $d=1,2$.
carries on to any random walk in a balanced, elliptic environment, therefore answering Conjecture 1.2 in the affirmative. Recall that a random walk is transient or recurrent depending on $G^{\mathscr{E}({\omega})}(x,x)<\infty$ or $G^{\mathscr{E}({\omega})}(x,x)=\infty$.
To state the next result, which connects balanced and general unbalanced regimes, we need the notion of embedded environments. The motivation for the construction can be understood only after the reader gets acquainted with the control argument which will be demonstrated accordingly.
3.1. Embedded Environments
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$.
The embedded environment function, denoted by $\Lambda\equiv\Lambda_{\omega}$, is a balanced environment function given by
$\Lambda_{\omega}(x,v)$ $=\frac{\omega(x,v)+\omega(x,-v)}{2}$ (1)
for every $v\in V_{d}$.
We shall assume without loss of generality that the environment function is Identity.
Theorem 3.7.
Assume $d\geq 1$. Let $\mu$ be an stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. Then suppose the environment process for the Embedded Environment satisfies the conclusions of 3.1.
Then there exists a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 1}$ is stationary and ergodic.
This theorem, combined with Theorem 3.1, gives us the following that
Theorem 3.8.
Let $d\geq 1$ and let $\mu$ a stationary ergodic environment law on $\mathcal{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
  1. (i)
    There exists a deterministic vector $v\in\mathbb{R}^{d}$ such that
    $\lim_{n\to\infty}\frac{X_{\omega}(n)}{n}=v\quad\text{$\omega-\mu^{0}$-a.e}.$
  2. (ii)
    $\frac{X_{\omega}([nt])-\sum_{i=0}^{[nt]}d(0,\tau_{X_{\omega}(m)}\omega)}{\sqrt{n}}$
    converges to the standard Brownian Motion on $\mathbb{R}^{d}$ with non-degenerate covariance matrix $\omega$-$\mu$ a.e.
3.2. Remarks:
The embedding function allows us to couple two random walks in $\mathbb{Z}^{d}$ namely the original and a reflected walk. This is motivated by the fact that the averaged random walk is balanced. Through the control argument, we pass on from the balanced to the general case. The Action of Linear Transformations proves crucial in performing this step. This might and should appear strange without prior context, but we shall see this arising naturally as an artifact of our proof technique.
4. An outline of the Proofs
Let us begin this section by seeing why an invariant measure for the local process guarantees answers to Conjectures 1.1. We assume for now that Theorems 3.1 and 3.7 to be true and prove 3.8
4.1. Proof of Theorem 3.8:
First note that by Theorems 3.1 and 3.7 we have a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 1}$ is stationary and ergodic since Assumptions 2.2 and 2.3 are satisfied. Call this measure $\mathbb{Q}$. Assume without loss of generality that $X_{\omega}(0)=0$.
Proof of Conjecture 1.1: We work with the environment following the distribution $\mathbb{Q}$. Consider the filtrations $\mathcal{F}_{k}=\sigma(X_{\omega}(0),\dots,X_{\omega}({k-1}))$ for $k\geq 1$. Then we can write the following expression,
$\frac{X_{\omega}(n)}{n}=\sum_{k=0}^{n-1}\frac{X_{\omega}(k+1)-X_{\omega}(k)-d(0,\tau_{X_{\omega}(k)}\omega)}{n}+\sum_{k=0}^{n-1}\frac{d(0,\tau_{X_{\omega}(k)}\omega)}{n}.$
Observe under the filtration sequence $\{\mathcal{F}_{n}\}_{n\geq 1}$
$\sum_{k=0}^{n-1}\frac{X_{\omega}(k+1)-X_{\omega}(k)-d(0,\tau_{X_{\omega}(k)}\omega)}{n},$
is a martingale.
By the Strong Law of Large Numbers for bounded martingale difference sequences,
$\sum_{k=0}^{n-1}\frac{X_{\omega}(k+1)-X_{\omega}(k)-d(0,\tau_{X_{\omega}(k)}\omega)}{n}\to 0\hskip 5.0pt\mathbb{P}_{0}^{\omega}\ \ \text{a.e.}$
And by Birkhoff’s Ergodic Theorem [2],
$\sum_{k=0}^{n-1}\frac{d(0,\tau_{X_{\omega}(k)}\omega)}{n}\to\int_{\mathcal{S}(d)}d(0,{\omega})d\mathbb{Q}\hskip 5.0pt\ \ \mathbb{Q}\ \ \text{a.e.}$
Combining both gives us the theorem $\mathbb{Q}_{0}=\mathbb{Q}\rtimes\mathbb{P}_{0}^{\omega}$ a.e., which by mutual absolute continuity of measures shows it holds $\mu^{0}=\mu\rtimes\mathbb{P}_{0}^{\omega}$ a.e.
Proof of Theorem 3.8(ii): First we note that under the sequence of filtrations $\{\mathcal{F}_{k}\}_{k\geq 0}$,
$U_{k}=X_{\omega}(k)-\sum_{m=0}^{k}d(0,\tau_{X_{\omega}(m)}\omega)$
is a martingale with bounded increments. Denote
$T(k)=\omega(X_{\omega}(k)).$
Therefore
$\Sigma_{k}=cov[X_{\omega}(k)-X_{\omega}(k-1)-d(0,\tau_{X_{\omega}(k-1)}\omega)|\mathcal{F}_{k}]$
has the entries
$(\Sigma_{k})_{ij}=\begin{cases}T(k)_{i}+T(k)_{i+d}-(T_{i}(k)-T_{i+d}(k))^{2}&\text{if }i=j,\\ -[T_{i}(k)-T_{i+d}(k)][T_{j}(k)-T_{j+d}(k)],&\text{if }i\neq j\end{cases}$
for $1\leq i,j\leq d$.
Therefore by 3.7,
$\frac{1}{n}\sum_{k=0}^{n-1}\Sigma_{k}\to\Sigma\quad\omega\penalty 10000\ \mu\textit{-a.e.}$
where $\Sigma$ is a positive definite matrix under assumption 2.3. Therefore, we conclude the theorem via the invariance principle for Martingales.
The proof of Theorem 3.5 is postponed and will be shown as a natural consequence of our uniqueness principle. Thus, we shall devote the rest of the paper to proving Theorems 3.1 and Theorem 3.7. We divide the proof into sections. Although each section may be read independently without loss of continuity, we recommend following the order presented, as it provides the intended motivation.
4.2. Discussion on Theorems 3.1, Corollary 3.2 and Theorem 3.5:
This section provides an alternate proof of Theorem 3.3 but generalises it to the general case of stationary, ergodic environments and arbitrary environment functions. This theorem is important for verifying the conditions of Theorem 3.7 under the assumptions of Theorem 3.8.
The two main themes in this proof are the use of Periodised Environments as used in [7] and our uniqueness principle. We shall prepare the reader with the prerequisites for the proof in the next section and fully motivate the proof. The proof in [7] adopts the arguments of Papanicolaou and Varadhan [8] for diffusion processes with random coefficients for the random walk for uniformly elliptic balanced environments. Our main step is to refine the arguments in [7] for non-uniformly elliptic balanced environments under assumptions 2.2 and 2.3 to get our uniqueness principle. The ideas of Krylov [6] were used to estimate the solutions of the discrete Monge-Ampere Equation by Lawler. We adopt these ideas to bound a certain resolvent as in [7] under assumptions 2.2 and 2.3. We further use this uniqueness principle to derive Theorem 3.5. We emphasise that Theorem 3.5 is not a result on Random Walk in Random Environments but on general balanced elliptic walks.
4.3. Discussion on Theorems 3.7:
There are two core themes in this proof. The first being the use of Action of Linear Transformations on the Environment. This approach is novel to the best of my knowledge. Let us describe this theory in brief. When we talk of shifts on the environment $\tau_{x}\omega$, we recall it is given by the relation
$\tau_{x}\omega(y)=\omega(x+y)\quad\forall x,y\in\mathbb{Z}^{d}.$
Through the Action theory, we talk of how general linear transformations on $\mathbb{Z}^{d}$ behave on the environment. Precisely speaking, we study expressions of the form $\tau_{Tx}\omega$ or $\tau_{x}T.\omega$, where $T\in\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}^{d},\mathbb{Z}^{d})$. These notions prove extremely fundamental to our arguments as we shall see.
We shall prepare the reader accordingly in the next section for the proof.
The second core theme in this proof is to pass from the embedded environment to the original environment by exhibiting stationarity of the local process through a control argument using Bernoulli Random Variables Independent of the environment. This allows us to “couple” the original walk and a reflected walk. The control argument allows us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process. The construction is specific to the proof and shall be demonstrated accordingly during the proof.
The main hurdle in the proof is to show stationarity of the local process. The ergodicity of the local process follows by adapting arguments from [3] suitably, which shall be demonstrated.
5. Preparation for the Proof of Theorem 3.1, Corollary 3.2 and Theorem 3.5 :
Fix an environment ${\omega}\in\mathcal{S}(d)$ and $\mathscr{E}$ a balanced measurable environment function on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$.
We first state our uniqueness result. We shall use this result crucially in the upcoming sections.
5.1. Uniqueness Principle:
Define the following:
$D_{n}=\{x\in\mathbb{Z}^{d}:|x|_{1}\leq n\}$
$\partial D_{n}=\{x\in D:|x|_{1}=n\}$
and
$D_{n}^{\circ}=D\setminus\partial D_{n}.$
Let $x\in D_{n}^{\circ}$. Define the second-order difference operators for a concave function $z:D_{n}\to\mathbb{R}$ given by
$\Delta_{i}z(x)=z(x+e_{i})+z(x-e_{i})-2z(x)\quad i=1,....,d$
Define the hitting time
$\tau_{n}$ $=\inf\{m:X_{\mathscr{E}(\omega)}(m)\in\partial D_{n}\}.$
Let $f:D_{n}\to[0,\infty)$ be a function with $f\equiv 0$ on $\partial D_{n}$. Define the class of concave functions $\mathscr{A}(\mathscr{E},f)$ containing concave functions $u(x)$ satisfying the properties
  1. 1.
    $u\equiv 0$ on $\partial D_{n}$.
  2. 2.
    $L_{\mathscr{E}(\omega)}u(x)\leq-f(x)\quad\forall x\in D_{n}^{\circ}$.
Theorem 5.1 (Uniqueness Principle).
Let $(\omega,\mathscr{E})$ be any balanced, elliptic environment. There exists an unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that $L_{\mathscr{E}(\omega)}z(x)=-f(x)\quad\forall x\in D_{n}^{\circ}$ given by
$z(x)=\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)})(x)]$
where the expectation is taken assuming $X_{\mathscr{E}(\omega)}(0)=x\in D_{n}^{\circ}$. Furthermore, this function is also given alternately by
$z(x)=\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x),\quad\forall x\in\mathbb{Z}^{d}.$
Now we return to Theorem 3.1 and Corollary 3.2. Assume without loss of generality, $X_{\mathscr{E}(\omega)}(0)=0$.
To prove this corollary, one notes that the random walk under consideration is a martingale. Define
$Z_{j}=(Z_{j}^{1},\dots,Z_{j}^{d})=X_{\mathscr{E}(\omega)}(j)-X_{\mathscr{E}(\omega)}(j-1)\penalty 10000\ \text{for j}\geq 1.$
Define $\tau_{X_{\mathscr{E}(\omega)}(j)}\omega=Y_{j}$. Then
$\mathbb{P}[Z_{j}^{i}=e_{i}]=\mathbb{P}_{x}[Z_{j}^{i}=-e_{i}]=\mathscr{E}_{i}(Y_{j}(0))$
Under the filtration
$\mathcal{F}_{j}=\sigma(X_{\mathscr{E}(\omega)}(0),\dots,X_{\mathscr{E}(\omega)}(j-1))$
for $j\geq 1$. One can check that
$\mathbb{E}[Z_{j}|\mathcal{F}_{j}]=0.$
This shows our walk
$X_{\mathscr{E}(\omega)}(j)=\sum_{i=1}^{j}Z_{j}$
is a martingale. Furthermore, we have the relation
$cov(Z_{j}^{i_{1}},Z_{j}^{i_{2}}|\mathcal{F}_{j})=\begin{cases}0&\text{if $i_{1}\neq i_{2}$ }\\ 2\mathscr{E}_{i_{1}}(Y_{j-1})(0)&\text{if $i_{1}=i_{2}$}.\end{cases}$
Let
$V_{n}^{i}=\sum_{j=0}^{n-1}2\mathscr{E}_{i}(Y_{j}(0)).$
Then the invariance principle for martingale gives
$W_{n}(t)=(W_{n}^{1}(t),\dots,W_{n}^{d}(t))$
converges in distribution to the standard Brownian motion on $\mathbb{R}^{d}$ where $W_{n}^{i}$ is given by
$W_{n}^{i}(t)=\frac{\sum_{j=1}^{[nt]}Z_{j}}{\sqrt{V_{n}^{i}}}.$
Suppose there exists $b\in S(d)$ such that,
$\lim_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n-1}\mathscr{E}_{i}[\omega(X_{\mathscr{E}(\omega)}(k))]=b_{i}\quad i=1,2,\dots,d.$
Then, under the assumption 2.3, one can conclude the corollary. This is an ergodic theorem, and therefore one can conclude the invariance principle if we prove the following,
Theorem 5.2.
$\exists b\in S(d)$ such that
$\lim_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n-1}\mathscr{E}[\omega(X_{\mathscr{E}(\omega)}(k))]=b.$
for $\omega$-$\mu\penalty 10000\ \text{a.e.}$
We prove this by finding an appropriate measure mutually absolutely continuous to $\mu$ such that the ergodic theorem holds.
Recall that $Y_{j}$ is a Markov chain with the generator
$K^{\mathscr{E}}_{d}(g(Y_{j}))=\sum_{i=1}^{d}\mathscr{E}_{i}(Y_{j})(0)[g(\tau_{e_{i}}Y_{j})+g(\tau_{-e_{i}}Y_{j})]$
for all $g:\mathcal{S}(d)\to\mathbb{R}^{d}$. Therefore 3.2 is equivalent to the following theorem,
Theorem 5.3.
Let $\mu$ be a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. Then under assumptions 2.2 and 2.3, there exists a stationary ergodic measure $\lambda$ mutually absolutely continuous to $\mu$ under which $\{Y_{j}\}$ is stationary and ergodic.
To prove stationarity of the chain is enough to conclude ergodicity via arguments in [3]. Under 5.3,
$b=\int_{\mathcal{S}(d)}\mathscr{E(}\omega)d\lambda.$
We shall work towards proving Theorem 5.3 under assumptions 2.2 and 2.3. We shall now introduce the notion of Periodised Environments as mentioned in section 4.2, fundamental to the proof of 5.3.
5.2. Periodised Environments
Define the equivalence relation $\sim$ by
$(x_{1},x_{2},\dots,x_{d})\sim(y_{1},y_{2},\dots,y_{d})\iff\frac{(x_{i}-y_{i})}{2n}\in\mathbb{Z}\penalty 10000\ \text{for each}\penalty 10000\ i=1,2,..d.$
Let $\mathbb{Z}^{d}/\sim=T_{n}$. Clearly $|T_{n}|=(2n)^{d}$. We shall also identify $T_{n}$ as the set in $\mathbb{Z}^{d}$ given by
$T_{n}=\{-n+1,-n+2,\dots,0,1,2,\dots,n\}^{d}.$
Definition 5.4 (Periodic Environment).
We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$.
[…]
$\hat{\omega}(x)$ $=\omega([x]_{\sim}),$
$\mathscr{E}(\hat{\omega})(x)$ $=\mathscr{E}(\omega)([x]_{\sim})$
where $x\in\mathbb{Z}^{d}$. Let $\mathcal{S}_{n}$ denote the set of all such $\hat{\omega}$. Note that $\mathcal{S}_{n}\subset\mathcal{S}(d)$.
The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the generator $L_{\mathscr{E}(\hat{\omega})}$.
For a function $g:T_{n}\to\mathbb{R}$, we extend it over $\mathbb{R}^{d}$ as
$g(x)=g([x]_{\sim}).$
Denote by $R_{n}^{\omega}$ the resolvent given by
$R_{n}^{\omega}g(x)=\sum_{j=0}^{\infty}(1-\frac{1}{n^{2}})^{j}L_{\mathscr{E}({\omega})}^{j}g(x)$
If $g:T_{n}\to\mathbb{R}$. Define the $l_{p}$($1\leq p\leq\infty)$ norms with respect to the normalized counting measure on $T_{n}$ by
[…]
As in the original paper by [7], in the following lemma we state a bound on the resolvent which in turn proves 5.3.
Lemma 5.5.
There exists $c_{1}\equiv c_{1}(d)>0$ such that $\forall\omega\in\mathcal{S}_{n}$,
[…]
To show that 5.5 $\implies$ 5.3, we first state a lemma by Parthsarathy[9] under the assumption that $\mu$ is stationary.
Lemma 5.6.
[Lemma $5$ of [7]] For each $n$ there exists $\omega_{n}\in\mathcal{S}_{n}$ such that the sequence of measures $\mu_{n}$ given by
$\mu_{n}=\frac{1}{(2n)^{d}}\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}$
converges weakly to $\mu$.
First we see the proof of 5.3, assuming $5.$
Proof of Theorem 5.3: Let $\{\omega_{n}\}_{n\geq 1}$ be the sequence satisfying 5.6. Consider $R_{n}=R^{\omega_{n}}_{n}$ as the following map between normed spaces,
$R_{n}:l^{d}(T_{n})\to l^{\infty}(T_{n}).$
Then the dual of $R_{n}$ is the map
$R^{*}_{n}:l^{1}(T_{n})\to l^{\frac{d}{d-1}}(T_{n})$
Let the invariant measure for the random walk on the group $T_{n}$ given by $X_{\mathscr{E}(\omega_{n})}(k)]_{\sim}$ be $\frac{\phi_{n}(x)}{(2n)^{d}}$. Note that $\phi_{n}$ is the density with respect to the normalised counting measure on $T_{n}$. Therefore, one can check (or refer to Section 2 of [5]) the invariant measure for $\{\omega_{n}(\hat{X}_{k})\}_{k\geq 0}$ is given by
$\lambda_{n}(x)=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
Since $\mathcal{S}(d)$ is compact with respect to the product topology, there exists a subsequence $n_{k}$ such that
$\lambda_{n_{k}}\to\lambda\penalty 10000\ \text{weakly}$
for some measure $\lambda$ on $\mathcal{S}(d)$ and $K^{\mathscr{E}}_{d}$ is invariant under $\lambda$.
Now observe that
[…]
By log-convexity of $l_{p}$ norms and Holder’s Inequality,
[…]
Therefore
[…]
Since
[…]
This shows
$\frac{d\lambda_{n}}{d\mu_{n}}||_{\frac{p}{p-1}}<\infty\quad\mu\penalty 10000\ \text{-a.e.}$
Since $\lambda_{n}$ extends as a measurable function over the entire $\mathbb{Z}^{d}$, we have that the family of functions
$\big\{\frac{d\lambda_{n}}{d\mu_{n}}\big\}_{n\geq 1}$
is uniformly integrable with respect to $\mu$. Therefore by Dunford-Pettis Theorem, $\lambda\ll\mu$ and
$\int_{\mathcal{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<\infty$
Referring to arguments from [3], $\lambda$ is ergodic and $\lambda\ll\mu$ and $\lambda\gg\mu$. This concludes the proof.
Thus, in view of the arguments in this section, we shall prove Lemma 5.5 in the next section.
6. Proof of Lemma 5.5:
We assume the Uniqueness Principle[Theorem 5.1] to be true. We first deduce Lemma 5.5 under this assumption and then prove Theorem 5.1 in the next section
Proof of Lemma 5.5 : Let $\omega\in\mathcal{S}_{n}$ and let $f:D_{n}\to[0,\infty)$. Let
$\tau_{n}$ $=\inf\{m:X_{\mathscr{E}(\omega)}(m)\in\partial D_{n}\}$
$Qf(x)$ $=\mathbb{E}_{x}[\sum_{k=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(k))]$
where the expectation is taken assuming $X_{{\mathscr{E}(\omega)}}(0)=x\in D_{n}^{\circ}$
The first part of this proof is dedicated to proving the following lemma,
Lemma 6.1.
For every $f:D_{n}\to[0,\infty),$
[…]
for some $c_{2}\equiv c_{2}(d)>1$.
Lemma 6.2 is equivalent to the Lemma 6.1 which states that,
Lemma 6.2.
Let $f:D_{n}\to[0,\infty)$ be any function. Let $z$ be the concave function guaranteed by Theorem 5.1. Then there exists $c_{3}>1$, dependent only on the dimension such that
[…]
To see the equivalence one notes the fact that $z(x)=Qf$.
Proof of Lemma 6.2: To prove the Lemma, we consider the Monge-Ampere Operator for a given concave function $z$ given by
$Mz(x)=\prod_{i=1}^{d}\Delta_{i}z(x)$
Note that from Theorem 5.1 one has that
$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)=f(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x)$
Define the sets for $x\in D_{n}^{\circ}$
$I(x)=\{a\in\mathbb{R}^{d}:z(x+e_{i})-z(x)\leq a_{i}\leq z(x)-z(x-e_{i}),\text{for}\penalty 10000\ i=1,2,\dots,d\}.$
Observe that
$meas(I(x))\leq\frac{f(x)^{d}}{c(\mathscr{E},x)^{d}},$
where $meas(\cdot)$ is the Lebesgue measure on $\mathbb{R}^{d}$.
Now we state an easily provable lemma as in [7].
Lemma 6.3 (Lemma 8 of [7]).
Let $a\in\mathbb{R}^{d}.$ and let $r(x)=a.x+b$ for $b>0$. If $r(x)\geq z(x)$ for every $x\in D_{n}$ and $\exists x_{0}\in D_{n}^{\circ}$ such that $r(x_{0})=z(x_{0})$, then $a\in I(x_{0})$.
Define the set
[…]
Therefore there exists an $b$ such that $a.x+b\geq z(x)$ and $\exists x_{0}$ such that $a.x_{0}+b=z(x_{0})$. Observe that
[…]
Therefore
$A\subset\bigcup_{x\in D_{n}^{\circ}}I(x),$
which gives that
[…]
where $c_{4}$ depends on $d$. Therefore,
[…]
Choose $c_{3}=c_{4}\vee 1$. This completes the proof of 6.2(c).
Now to bound the resolvent[recall the statement of Lemma 5.5], first we take $m=Kn$ such that $T_{n}\subset D_{m}$. We shall specify this $K$. Assume for now that this $K$ only depends on $d$, and it is a natural number.
We now bound the resolvent for non-negative $g:T_{n}\to\mathbb{R}$(and extended as mentioned before) as follows,
$\mathbb{E}_{x}\big[\sum_{j=0}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega)}(j)\big)]\leq\mathbb{E}_{x}\big[\sum_{j=0}^{\tau_{Kn}}g(X_{\mathscr{E}(\omega)}(j))\big]+\mathbb{E}_{x}\big[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega)}(j)\big)],$
where $x\in T_{n}$.
Using 6.1, one has that
[…]
where $c^{\prime}\equiv c^{\prime}(d)$.
The last inequality follows due to the periodic nature of $g$.
For the other summand, we write it as follows,
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\Bigg]$ $=\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}>n^{2}\}}\Bigg]$ (2)
$\quad+\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg].$
The first sum can be bounded as follows,
[…]
For the other summand, observe that
$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg].\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\penalty 10000\ ||R_{n}g||_{\infty}.$
The last line follows from the Strong Markov property as
$\{\tau_{Kn}\leq n^{2}\}\in\mathcal{F}_{\tau_{Kn}}=\{A:\forall\,t\in\mathbb{N},\ \{\tau_{Kn}\leq t\}\cap A\in\mathcal{F}_{t+1}\}$
$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right)$
Since $|X_{\mathscr{E}(\omega)}(n)|_{1}^{2}$ is a submartingale, we have by Doob’s inequality that
$\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]=\mathbb{P}_{x}[\max_{0\leq m\leq n^{2}}|X_{{\mathscr{E}(\omega)}}(m)|_{1}^{2}\geq K^{2}n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}.$
Further more observe that since $x\in T_{n}$ we must have $|x|^{2}_{1}\leq n^{2}d^{2}$ for every $x\in T_{n}$. Therefore, we have
$\sup_{x\in T_{n}}\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}\leq\frac{d^{2}+1}{K^{2}}.$
Now choose $K$ such that
$\frac{d^{2}+1}{K^{2}}$ $<\frac{1}{2}-\frac{1}{e},\quad T_{n}\subset D_{m}$
Therefore, combining the deductions we have that
[…]
By construction, the assumptions of Lemma 5.5 are satisfied, thereby the proof.
Therefore, it is now enough to prove Theorem 5.1. We shall in the next section also see how this leads to a proof of Theorem 3.5.
7. Proof of Theorem 5.1 and Theorem 3.5
Proof of Theorem 5.1: We begin by noting that the class $\mathscr{A}(\mathscr{E},f)$ is non-empty. Indeed consider
$u(x)=n(n+1)-|x|_{1}(|x|_{1}+1).$
Observe that since $\Delta_{i}u(x)\leq-2$, using ellipticity, one can conclude that $\exists\gamma>0$ such that $\gamma u\in\mathscr{A}$.
We now make two easy observations without proof. The first one is that
$z_{1},z_{2}\in\mathscr{A}(\mathscr{E},f)\implies z_{1}\wedge z_{2}\in\mathscr{A}(\mathscr{E},f).$
The second observation is that
$z(x)=\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x)\in\mathscr{A}(\mathscr{E},f).$
We claim that this function satisfies the second condition of the Theorem. Indeed if it does not then $\exists x_{0}\in D_{n}^{\circ}$ such that
$L_{\mathscr{E}(\omega)}u(x_{0})<-f(x_{0}).$
One can choose $0<\beta<f(x_{0})$ such that
$\sum_{i=1}^{d}\mathscr{E}(\omega)(x_{0},e_{i})[z(x_{0}+e_{i})+z(x_{0}-e_{i})-2\beta]<-f(x_{0}).$
It is again easy to show that
$z^{\prime}(x)=\begin{cases}\beta,&\text{if }x=x_{0},\\ z(x),&\text{otherwise}.\end{cases}$
is a member of $\mathscr{A}(\mathscr{E},f)$. This contradicts the choice of $z(x)$.
One also notes that for every $j\geq 1$,
$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]=\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\omega}(k))$
Letting $k\to\infty$ completes the proof.
Proof of Theorem 3.5: Fix any balanced elliptic environment $(\omega,\mathscr{E})$. Let $f:D_{n}\to[0,\infty)$ be a function. Observe that for every concave function $u\in\mathscr{A}(\mathscr{E},f)$ one has
$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\leq\frac{c(\mathscr{E},x)^{d}}{2d}\sum_{i=1}^{d}\Delta_{i}u(x)$
Denote by the $i$ the environment function which gives $\frac{1}{2d}$ on all the coordinates except the first one where it assigns 0.
Therfore we have that
$\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)\subset\mathscr{A}(\mathscr{E},f)$
Denote by $\{X_{j}\}_{j\geq 0}$ to be the simple symmetric random walk. Then we have by Theorem 5.1 that
$\inf_{u\in\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)}u(x)\geq\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x)$
or by using the other form,
$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}]\geq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))].$
Similarly we note that
$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\geq\sum_{i=1}^{d}\Delta_{i}u(x).$
This shows that
$\mathscr{A}(\mathscr{E},f)\subset\mathscr{A}\left(i,\frac{f(x)}{2d}\right)$
As above, one can show that
$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}f(X_{j})]$
Therefore, we have established the following result.
Theorem 7.1.
Let $f:D_{n}\to[0,\infty)$ be a function. Let $(\omega,\mathscr{E})$ be a balanced elliptic environment. Then for all $x\in D_{n}^{\circ}$, we have the inequality,
$\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}]$
One can now conclude Theorem 3.5 by taking the function $f:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$
$f(u)=\boldsymbol{1}_{x}(u)$
and letting $n\to\infty$.
8. Preparation for the proof of Theorem 3.7:
We begin this section by discussing the theory of Action of Linear Transformations. As mentioned earlier, for this proof we shall assume $\mathscr{E}=\operatorname{Id}$.
8.1. Action of Linear Transformations:
Fix $\omega\in\mathcal{S}(d)$. For this proof For a given linear transformation $T\in\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}^{d},\mathbb{Z}^{d})$, the action of $T$ on $\omega=\{\omega(x,v)\}_{x\in\mathbb{Z}^{d},v\in V_{d}}$ is given by
$T.\omega(x,v):=T(\omega)(x,v)=\omega(Tx,v).$
Similarly, we also derive the form of the combined action of the shift and the transformation, which we state in the following theorem.
Theorem 8.1.
$\tau_{x}T.\omega=T(\tau_{Tx}\omega)$
Proof.
Observe
$(\tau_{x}T.\omega)(y,v)=T.\omega(x+y,v)=\omega(Tx+Ty,v).$
We also have that
$\omega(Tx+Ty,v)=T(\tau_{Tx}\omega)(y,v).$
This establishes the assertion. We also describe the action on the transition kernel. We note that we derive these for permutation maps, i.e.,
$T(V_{d})=V_{d}.$
It is straightforward that such a map is invertible.
The induced random walk under $T$ in $\mathbb{Z}^{d}$ started from $x\in\mathbb{Z}^{d}$ is given by
$\mathbb{P}_{0}^{\omega}(T(X_{n})=T(X_{n-1})+T(e_{i}))=\mathbb{P}_{0}^{\omega}(X_{n}=X_{n-1}+e_{i})=\omega(X_{n-1},e_{i})$
Call the transition kernel of the environment under this new random walk, i.e., the process $\tau_{T(X_{n})}\omega$, $K^{T}_{d}(\omega,B)$. Then we have the result,
Theorem 8.2.
$K_{d}^{\operatorname{Id}}\equiv K_{d}(T.\omega,T(B))=K^{T}_{d}(\omega,B)$ for every $B\in{\mathcal{B}(d)}$
Proof.
Observe that,
$K_{d}^{T}(\omega,B)=\sum_{i=1}^{d}\Big[\omega(o,e_{i})\,\delta_{\tau_{T(e_{i})}\omega}(B)+\omega(o,-e_{i})\,\delta_{\tau_{T(-e_{i})}\omega}(B)\Big]+\omega(o,o)\delta_{\omega}(B).$
Note that $T.\omega(o,e)=\omega(o,e)$ and by Theorem 8.1,
$\{\tau_{e}{T.\omega}\in T(B)\}=\{T(\tau_{Tx}\omega)\in T(B)\}=\{\tau_{Tx}(B)\in B\}\quad\penalty 10000\ \forall e\in V_{d}.$
This completes the proof. ∎
This also shows another fact, that is,
Corollary 8.3.
If $\tau_{X_{\mathscr{E}(\omega)}(n)}\omega$ has a stationary measure $Q$, $\tau_{T(X_{\mathscr{E}(\omega)}(n))}\omega$ is also stationary under $Q\circ T$.
Proof.
$\int_{\mathcal{S}(d)}K_{d}^{T}(\omega,B)dQ\circ T=\int_{\mathcal{S}(d)}K_{d}(T.\omega,T(B))dQ\circ T=Q\circ T(B)\quad\forall B\in\mathcal{B}({d})^{\prime}$
We shall use these notions fundamentally in our proof. We also note that showing stationarity of the local process is sufficient to exhibit its ergodicity, which we state in the following lemma. We shall prove the lemma by adapting arguments from [3].
Lemma 8.4.
If $\mathbb{Q}$ is a measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$ , such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 1}$ is stationary then it is also ergodic.
Let $\mathcal{S}^{\prime}=S(d)^{\mathbb{N}}$ denote all $S(d)$ valued trajectories of the local process. Let $T$ denote the canonical shifts or the $n^{th}$ iterate of the process. Denote by $\tilde{\mathcal{B}}$ the product $\sigma$-algebra. Let $A$ be a set which is invariant under $T$, i.e., $T^{-1}A=A$.
Let $h(\omega)=\tilde{P}_{\omega}(A)$ where $\tilde{P_{\omega}}$ is the law if the chain starts from $\{\omega(x)\}_{x\in\mathbb{Z}^{d}}\sim\mathbb{Q}$. We must show
$\tilde{P}_{\mathbb{Q}}(A)=\int_{\mathcal{S}(d)}\tilde{P}_{\omega}(A)d\mathbb{Q}(\omega)=0\ \ or\ \ 1.$
Denote $\bar{\omega}(n)=\tau_{X_{\omega}(n)}\omega$.
Start by noting that under the canonical filtrations sequence $\mathscr{G}_{k}=\sigma(\omega,\bar{\omega}(1),\dots,\bar{\omega}(k-1))$,
$h(\bar{\omega}(n))$ is a $\tilde{P}_{\mathbb{Q}}$ martingale since
$\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}\circ T^{n}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=h(\bar{\omega}(n))\quad\mathbb{Q}\penalty 10000\ \text{-a.e.}$
By the Levy Upwards Theorem,
$h(\bar{\omega}(n))\to\boldsymbol{1}_{A}\qquad(n\to\infty)\quad\mathbb{Q}\penalty 10000\ \text{-\text{a.e.}}$ (3)
Next we show $h=\boldsymbol{1}_{B}$($\mathbb{Q}$ a.e.) for some $B\in\mathcal{B}(d)^{\prime}$. If not then $\mathbb{Q}(h\in[a,b])>0$ for some $0<a<b<1$.
But by Birkhoff’s Ergodic Theorem,
$\frac{1}{n}\sum_{k=0}^{n-1}\boldsymbol{1}\{h(\bar{\omega}(k))\in[a,b]\}\to\Phi=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}\{h(\omega)\in[a,b]\}|\mathcal{I}]\quad\mathbb{Q}\penalty 10000\ \text{-a.e.}$
where $\mathcal{I}$ is the $\sigma$-algebra of invariant events.
Observe,
$\mathbb{E}_{\mathbb{Q}}[\Phi]=\mathbb{Q}[h\in[a,b]]>0.$
But this contradicts 3 since $\Phi=0$ ($\mathbb{Q}$ a.e.) Now we note that from the martingale property and the Markov property,
$\boldsymbol{1}_{B}(\omega)=\mathbb{E}_{\mathbb{Q}}\Big[\,\boldsymbol{1}_{B}(\bar{\omega}(1))\,\big|\,\omega\,\Big]=K_{d}\,\boldsymbol{1}_{B}(\omega)\quad\mathbb{Q}\penalty 10000\ \text{-a.e.}$ (4)
Combining 4 with ellipticity,
$\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{e}\quad(\mu\text{-\text{a.e.} or }\mathbb{Q}\penalty 10000\ \text{-\text{a.e.}})\quad\forall\,e\in{V}_{d}.$ (5)
Iterating 5, we obtain
$\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{x}\quad(\mu\text{-\text{-a.e.} or }\mathbb{Q}\penalty 10000\ \text{-\text{a.e.}})\quad\forall\,x\in\mathbb{Z}^{d}.$
Since $\mu[\omega\in B]=\mu[\tau_{x}\omega\in B]$ for every $x\in\mathbb{Z}^{d}$,
$\boldsymbol{1}_{B}=\boldsymbol{1}_{B}\circ\tau_{x}\quad\mu\penalty 10000\ \text{\text{-a.e.} or}\penalty 10000\ \mathbb{Q}\penalty 10000\ \text{-\text{a.e.}}$ (6)
Consider the set
$\tilde{B}=\cap_{x\in\mathbb{Z}^{d}}\tau_{-x}B.$
Observe $(\tau_{x})^{-1}\tilde{B}=\tilde{B}$ which shows $\tilde{B}$ is invariant under shifts.
From 6 we get that $\mu(\tilde{B})=\mu(B)=0$ or $1$ (by ergodicity). From the mutual absolute continuity of $\mathbb{Q}$ and $\mu$, $\mathbb{Q}(B)=\tilde{P}_{\mathbb{Q}}(A)=0$ or $1$. This proves ergodicity of the local process in the environment $\mathbb{Q}$.
Thus, in view of the previous proof, we have reduced the proof to the following lemma.
Lemma 8.5.
Theorem 3.7 holds under its assumptions if the local process is stationary under some measure $\mathbb{Q}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ which is mutually absolutely continuous to $\mu$.
We shall develop in the next section the proof of Lemma 8.5.
9. Proof of Lemma 8.5:
The first part of the proof is to construct $\mathbb{Q}$.
Let $\lambda^{\prime}$ be an invariant measure of $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 1}$. The first make an observation on this measure.
Let $\alpha:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ given by $\alpha(e_{i})=-e_{i}\quad i=1,2,...,d$.
Lemma 9.1.
$\lambda^{\prime}$ is an $\alpha$-invariant measure. More precisely,
$\lambda^{\prime}\circ\alpha(B)=\lambda^{\prime}(B)\quad\forall B\in\mathcal{B}(d)^{\prime}$
Proof.
Start by noting that for every $B\in\mathcal{B}(d)^{\prime}$,
$\int_{\mathcal{S}(d)}K_{d}^{\Lambda}(\omega,B)d\lambda^{\prime}\circ\alpha=\int_{\mathcal{S}(d)}K_{d}^{\Lambda}(\omega,\alpha(B))d\lambda^{\prime}=\lambda^{\prime}\circ\alpha(B).$
This shows $\lambda^{\prime}\circ\alpha$ is also an invariant stationary measure for the environment process. Furthermore $\lambda^{\prime}\circ\alpha$ is mutually absolutely continuous to $\mu\circ\alpha$. One can also show that $\mu\circ\alpha$ is stationary ergodic with respect to the group-of-shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$.
Let $g:\mathcal{S}(d)\to\mathbb{R}$ be a measurable function. By Birkhoff’s Ergodic Theorem,
$\frac{1}{n}\sum_{i=0}^{n-1}g(\tau_{X_{\Lambda(\omega)}(n)}\omega)$ $\to\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\quad\omega\penalty 10000\ \mu\textit{-a.e.}$ (7)
$\frac{1}{n}\sum_{i=0}^{n-1}g(\tau_{X_{\Lambda(\omega)}(n)}\omega)$ $\to\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\circ\alpha\quad\omega\penalty 10000\ \mu\circ\alpha\textit{-a.e.}$ (8)
Therefore, we conclude that,
$\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}=\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\circ\alpha,$
which completes the proof. ∎
Fix $\omega\in\mathcal{S}(d)$.
Recall that $\{X_{\omega}(n)\}_{n\geq 1}$ is the random walk with the law
$\mathbb{P}[X_{\omega}(1)=x+v\Big|X_{\omega}(0)=x]=\omega(x,v)\quad v\in V_{d}$ (9)
The environment process is therefore given by $K_{d}^{\operatorname{Id}}\equiv K_{d}$.
Define the reflected walk, $\{X_{\alpha(\omega)}(n)\}_{n\geq 1}$ with the law
$\mathbb{P}[X_{\alpha(\omega)}(1)=x+v\Big|X_{\alpha(\omega)}(0)=x]=\omega(x,-v).$ (10)
The environment process is given by
$G_{d}(\omega,B):=K^{\alpha}_{d}(\omega,B)=K_{d}(\alpha.\omega,\alpha.B)\quad\forall B\in\mathcal{B}(d)^{\prime}.$
We redefine $X_{\Lambda(\omega)}(n)$ in the following manner.
$L_{\Lambda(\omega)}g(X_{\Lambda(\omega)}(n))=\mathbb{E}[\gamma_{n}L_{\omega}g(X_{\Lambda(\omega)}(n))+(1-\gamma_{1})L_{\alpha(\omega)}g(X_{\Lambda(\omega)}(n))]\quad\forall g:\mathbb{Z}^{d}\to\mathbb{R},\penalty 10000\ \gamma_{n}\sim\text{Ber}\left(\frac{1}{2}\right)$
where $L_{\Lambda(\omega)},L_{\omega},L_{\alpha(\omega)}$ are the generators of $X_{\Lambda(\omega)},X_{\omega},X_{\alpha(\omega)}$ respectively and $\gamma_{n}$ is a Bernoulli Random Variable independent of the environment and position of the walk. Further $\{\gamma_{n}\}_{n\geq 1}$ is a sequence of i.i.d. Bernoulli random variables.
We therefore define the conditional environment process
$K_{d}^{\Lambda}(\omega,B)\Big|\gamma=K_{\gamma}(\omega,B)=\gamma K_{d}(\omega,B)+(1-\gamma)G_{d}(\omega,B).$
We now state an easy lemma without proof.
Lemma 9.2.
$K_{\gamma}(\alpha.\omega,B)=K_{1-\gamma}(\omega,\alpha.B)$
This is the control argument using Bernoulli variables as we had mentioned, and now we shall use it explicitly to demonstrate the stationarity of the local process.
We claim that $\mathbb{Q}=\lambda^{\prime}$.
Lemma 9.3.
Let $n\geq 1$. Let $B_{0},B_{1},...,B_{n}\in\mathcal{B}(d)^{\prime}$. Let the probability measure associated to $\omega$ be $\lambda^{\prime}$. Let $\bar{\omega}(n)=\tau_{X_{\Lambda(\omega)}(n)}\omega$. Then for every $1\leq k\leq n-1$, we have
$\lambda^{\prime}\Bigl[\omega\in B_{0},\,\bar{\omega}(1)\in B_{1},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,\gamma_{n}\Bigr]$
$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\ \bar{\omega}(k+1)\in\alpha(B_{k+1}),\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,1-\gamma_{k+1},\ldots,\gamma_{n}\Bigr]$ (11)
For $k=0$,
$\lambda^{\prime}\Big[\omega\in\alpha(B_{0}),\,\bar{\omega}(1)\in B_{1},\,\ldots,\,\bar{\omega}(k)\in B_{k},\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,\gamma_{n}\Bigr]$
$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\bar{\omega}(1)\in\alpha(B_{1})\ldots,\,\bar{\omega}(k)\in B_{k}\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,1-\gamma_{1},\ldots,\gamma_{k},\ldots,\gamma_{n}\Bigr]$ (12)
To see how this implies stationarity, take $B_{i}=\pi_{0}^{-1}(A_{i})$ where $A_{i}\in\mathcal{B}(d)$. Then
$\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{0}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})\Big|\gamma_{1},...,\gamma_{n}]=\frac{1}{2^{n}}\sum_{(\gamma_{1},...,\gamma_{n})\in\{0,1\}^{n}}\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{0}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})\Big|\gamma_{1},...,\gamma_{n}]$
The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{1}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$. Now conclude stationarity using the stationarity of $\bar{\omega}(n)$ and taking $\gamma_{i}=1$. It is clear that
$\tau_{X_{\Lambda(\omega)}(n)}\omega|\{\gamma_{i}=1\quad i=1,2,....,n\}=\tau_{X_{\omega}(n)}\omega.$
Therefore it is now enough to establish Lemma 9.3.
Proof of Lemma 9.3: We shall prove this result using induction. First for $n=1$, we note that
$\lambda^{\prime}[\alpha(\omega)\in B_{0},\bar{\omega}(1)\in B_{1}|\gamma]=\int_{\alpha(B_{0})}K_{\gamma}(\omega,B_{1})d\lambda^{\prime}(\omega)$
We perform the substitution $\omega\mapsto\alpha(\omega)$. The integral reduces to
$\int_{B_{0}}K_{\gamma}(\alpha.\omega,B_{1})d\lambda^{\prime}\circ\alpha$
By Lemma 9.1 and Lemma 9.2, the integral reduces to
$\int_{B_{0}}K_{1-\gamma}(\omega,\alpha.B_{1})d\lambda^{\prime}\circ\alpha=\lambda^{\prime}(\alpha(\omega)\in B_{0},\bar{\omega}(1)\in\alpha.B_{1}|1-\gamma)$
Suppose the result is true for $n=m$. Let $n=m+1$. Then we must verify the statement for $k=m+1$ only. For $k=m+1$, we write
$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times\alpha.B_{m}}K_{\gamma_{m+1}}(\omega_{m},B_{m+1})d\lambda^{\prime}(\omega_{m},...,\omega|\gamma_{1},...,\gamma_{m})$
Perform the substitution $(\omega_{m},...,\omega)\mapsto(\alpha.\omega_{m},...,\omega)$ to get
$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times B_{m}}K_{\gamma_{m+1}}(\alpha.\omega_{m},B_{m+1})d\lambda^{\prime}(\alpha.\omega_{m},...,\omega|\gamma_{1},...,\gamma_{m})$
Using Lemma 9.2, the integral reduces to
$\int_{B_{0}\times B_{1}\times B_{2}\times\cdots\times B_{m}}K_{1-\gamma_{m+1}}\bigl(\omega_{m},\alpha(B_{m+1})\bigr)\,d\lambda^{\prime}\bigl(\alpha(\omega_{m}),\ldots,\omega\,\Bigm|\,\gamma_{1},\ldots,\gamma_{m}\bigr)$
Now we write the above expression as
$\mathbb{E}\Big[\mathbb{E}\Big[\boldsymbol{1}_{\alpha.B_{m+1}}(\bar{\omega}(m+1))\Big|\bar{\omega}(m),1-\gamma_{m+1}\Big]\boldsymbol{1}_{\bar{\omega}(m)}(\alpha.B_{m})\boldsymbol{1}_{\bar{\omega}(m-1)}(B_{m-1})\dots,\boldsymbol{1}_{\omega}(B_{0})\Big|\gamma_{1},...,\gamma_{m}\Big]$
This reduces by definition to the desired form, which completes the proof.
Acknowledgements:
The author is deeply indebted to Dr. Soumendu Sundar Mukherjee of Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for helpful discussions and his immense help with the manuscript. The author also thanks his classmates Himadri Mandal and Aryama Ghosh for their helpful comments on the exposition and the manuscript in general.

### v2

@@
 We use a more general definition of environments using Environment Functions.
-As a corollary, we can deduce an invariance principle under these assumptions for balanced environments under some assumptions.
+As a corollary, we can deduce an invariance principle for balanced environments under the assumptions.
 We also use the uniqueness principle to show that any balanced, elliptic random walk must have the same transience behaviour as the simple symmetric random walk.
@@
 This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers.
-Roughly speaking, it says there is a deterministic limiting velocity of the walk, and the second question is about the probability of the walk escaping to infinity along a given direction.
-Both have been studied in detail in $d=1$ by Solomon[11] in 1975.
+Roughly speaking, it says there is a deterministic limiting velocity of the walk.
+Thid have been studied in detail in $d=1$ by Solomon[11] in 1975.
 In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
@@
 Varadhan in [12] showed that the sequence $X_{n}/n$ has at most two deterministic limit points $v_{1},v_{2}$ and if $v_{1}\neq v_{2}$ then there is a constant $a\geq 0$ such that $v_{1}=-av_{2}$.
-Furthermore, Nicholas Berger in [1] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0.
+Furthermore, Noam Berger in [1] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0.
 These results assume that the environment is uniformly elliptic and i.i.d..
@@
 We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest.
-Using the uniqueness we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks.
+Using the uniqueness, we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks.
 It is important to note that this section shall not deal with the notion of a Random Environment.
-This section is one of the most fundamental contributions, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
+This is one of the key contributions of this p, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
 Our other main contribution will be to deduce the ergodicity of a certain stochastic process, which we will call the local process.
@@
 This can be noted for Random Walks in Dirichlet Environments as studied by Christopher Sabot in [10].
-Our definition of environments allows us to transfer our results from the balanced case to the general environment, by coupling the random walk and a reflected walk.
-We then combine a novel technique which we shall name as the Action of Linear Transformations and another control technique to execute a proof.
+Our definition of environments allows us to transfer our results from the balanced case to the general environment by coupling the random walk and a reflected walk.
+We then combine a novel technique, which we shall name as the Action of Linear Transformations, and another control technique to execute a proof.
 As a consequence of our ergodic theorem, we shall also be able to deduce an invariance principle for general environments.
@@
 Our results hold for all uniformly elliptic RWRE, and the conditions we impose study the extent to which the condition of uniform ellipticity can be dropped.
+Acknowledgements:
+The author is an undergraduate student at the Indian Statistical Insitute, Kolkata.
+This work was completed during the third-year of his studies.
+He is deeply indebted to Dr.
+Soumendu Sundar Mukherjee of Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for immensely helpful discussions and his substantial help with the manuscript.
+The author also thanks his classmates Himadri Mandal and Aryama Ghosh for their helpful comments on the exposition and the manuscript in general.
 2. Prerequisites
@@
 Let $\mu$ be a probability measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. We call the environment function balanced if
-$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega(x))=\mathscr{E}_{i+d}(\omega(x))\quad i=1,2,...,d.$
+$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega(x))=\mathscr{E}_{i+d}(\omega(x))\quad i=1,2,\dots,d.$
 We shall call the environment elliptic if
@@
 $\mu[\tau_{-x}\omega\in A]=\mu[\omega\in A]\quad\forall A\in\mathcal{B}(d)^{\prime},\forall x\in\mathbb{Z}^{d}$
+$\forall B\in\mathcal{B}(d)^{\prime},\quad\mu\left[\cap_{x\in\mathbb{Z}^{d}}\tau_{x}B\right]\in\{0,1\}$
 As a special case, $\mu$ is said to be an independent and identically distributed (i.i.d.) law if $\{\omega(x)\}_{x\in\mathbb{Z}^{d}}$ is an i.i.d.
@@
 $\mathbb{P}[X_{\mathscr{E}(\omega)}(k+1)$ $=x+v|X_{\mathscr{E}(\omega)}(k)=x]=\mathscr{E}(x,v)\quad\forall v\in V_{d}$
-Therefore, the generators of this Markov Chain is given by
-$L_{\mathscr{E}(\omega)}(g(x))=\sum_{v\in V_{d}}\mathscr{E}(\omega)(x,v)(g(x+v)-g(x)),$
+Therefore, the kernel of this Markov Chain are given by
+$L_{\mathscr{E}(\omega)}(g(x))=\sum_{v\in V_{d}}\mathscr{E}(\omega)(x,v)g(x+v),$
 respectively for every $g:\mathbb{Z}^{d}\to\mathbb{R}$.
@@
 Assumption 2.2.
-Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d))$. Define the quantity $c(\mathscr{E},\omega)(x)$ as follows
-$c(\mathscr{E},\omega)(x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}$
+Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d))$. Define the quantity $c(\mathscr{E},x)$ as follows
+$c(\mathscr{E},x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}$
 Then we assume that
-$\int_{\mathcal{S}(d)}c^{-p}d\mu<\infty.$
+$\int_{\mathcal{S}(d)}c^{-p}d\mu<\infty,$
 for some $p>d$.
@@
 converges weakly to the standard Brownian Motion on $\mathbb{R}^{d}$ with non degenerate covariance matrix of the form $(b_{i}\delta_{ij})_{1\leq i,j\leq d}$ for $\omega$-$\mu$ a.e.
-These results notably weaken the assumptions of ellipticity and derives an alternate proof of Theorem 1 of [5] for $d\geq 2$, which states that
+These results notably weaken the assumptions of ellipticity and derive an alternate proof of Theorem 1 of [5] for $d\geq 2$, which states that
 Theorem 3.3 (Theorem 1(i) of [5]).
@@
 for some $p>d\geq 2$, then Corollary 3.2 holds.
-To prove the above results, we derive a uniqueness theorem for concave functions on $\mathbb{Z}^{d}$, which also gives the following theorem on the transcience of balanced elliptic walks.
+To prove the above results, we derive a uniqueness theorem for concave functions on $\mathbb{Z}^{d}$, which also gives the following theorem on the transience of balanced elliptic walks.
 The result compares the walk to the simple symmetric random walk on $\mathbb{Z}^{d}$ given by the law
@@
 To state the next result, which connects balanced and general unbalanced regimes, we need the notion of embedded environments.
-The motivation for the construction can be understood only after the reader gets acquainted with the control argument which will be demonstrated accordingly.
+The motivation for the construction can be understood only after the reader gets acquainted with the control argument, which will be demonstrated accordingly.
 3.1. Embedded Environments
@@
     There exists a deterministic vector $v\in\mathbb{R}^{d}$ such that
-    $\lim_{n\to\infty}\frac{X_{\omega}(n)}{n}=v\quad\text{$\omega-\mu^{0}$-a.e}.$
+    $\lim_{n\to\infty}\frac{X_{\omega}(n)}{n}=v\quad\omega\penalty 10000\ \text{$\mu$-a.e}.$
   2. (ii)
@@
 3.2. Remarks:
-The embedding function allows us to couple two random walks in $\mathbb{Z}^{d}$ namely the original and a reflected walk.
+The embedding function allows us to couple two random walks in $\mathbb{Z}^{d}$, namely the original and a reflected walk.
 This is motivated by the fact that the averaged random walk is balanced.
-Through the control argument, we pass on from the balanced to the general case.
+Through the control argument, we pass from the balanced to the general case.
 The Action of Linear Transformations proves crucial in performing this step.
@@
 is a martingale with bounded increments. Denote
-$T(k)=\omega(X_{\omega}(k)).$
+$T(k)=\omega(X_{\omega}(k-1)),\quad k\geq 1.$
 Therefore
@@
 Thus, we shall devote the rest of the paper to proving Theorems 3.1 and Theorem 3.7.
-We divide the proof into sections.
-Although each section may be read independently without loss of continuity, we recommend following the order presented, as it provides the intended motivation.
+We divide the proof into various sections in order to properly motivate the proof.
 4.2. Discussion on Theorems 3.1, Corollary 3.2 and Theorem 3.5:
@@
 Our main step is to refine the arguments in [7] for non-uniformly elliptic balanced environments under assumptions 2.2 and 2.3 to get our uniqueness principle.
-The ideas of Krylov [6] were used to estimate the solutions of the discrete Monge-Ampere Equation by Lawler.
+The ideas of Krylov [6] were used to estimate the solutions of the discrete Monge-Ampère Equation by Lawler.
 We adopt these ideas to bound a certain resolvent as in [7] under assumptions 2.2 and 2.3.
@@
 This allows us to “couple” the original walk and a reflected walk.
-The control argument allows us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process.
+This technique allows us to control the paths of the random walk which shall be key to the proof.
+The control argument will therefore allow us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process.
 The construction is specific to the proof and shall be demonstrated accordingly during the proof.
@@
 Let $x\in D_{n}^{\circ}$. Define the second-order difference operators for a concave function $z:D_{n}\to\mathbb{R}$ given by
-$\Delta_{i}z(x)=z(x+e_{i})+z(x-e_{i})-2z(x)\quad i=1,....,d$
+$\Delta_{i}z(x)=z(x+e_{i})+z(x-e_{i})-2z(x)\quad i=1,\dots,d$
 Define the hitting time
@@
 We prove this by finding an appropriate measure mutually absolutely continuous to $\mu$ such that the ergodic theorem holds.
-Recall that $Y_{j}$ is a Markov chain with the generator
+Recall that $Y_{j}$ is a Markov chain with the kernel
 $K^{\mathscr{E}}_{d}(g(Y_{j}))=\sum_{i=1}^{d}\mathscr{E}_{i}(Y_{j})(0)[g(\tau_{e_{i}}Y_{j})+g(\tau_{-e_{i}}Y_{j})]$
@@
 Define the equivalence relation $\sim$ by
-$(x_{1},x_{2},\dots,x_{d})\sim(y_{1},y_{2},\dots,y_{d})\iff\frac{(x_{i}-y_{i})}{2n}\in\mathbb{Z}\penalty 10000\ \text{for each}\penalty 10000\ i=1,2,..d.$
+$(x_{1},x_{2},\dots,x_{d})\sim(y_{1},y_{2},\dots,y_{d})\iff\frac{(x_{i}-y_{i})}{2n}\in\mathbb{Z}\penalty 10000\ \text{for each}\penalty 10000\ i=1,2,\dots,d.$
 Let $\mathbb{Z}^{d}/\sim=T_{n}$. Clearly $|T_{n}|=(2n)^{d}$. We shall also identify $T_{n}$ as the set in $\mathbb{Z}^{d}$ given by
@@
 where $x\in\mathbb{Z}^{d}$. Let $\mathcal{S}_{n}$ denote the set of all such $\hat{\omega}$. Note that $\mathcal{S}_{n}\subset\mathcal{S}(d)$.
-The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the generator $L_{\mathscr{E}(\hat{\omega})}$.
+The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the kernel $L_{\mathscr{E}(\hat{\omega})}$.
 For a function $g:T_{n}\to\mathbb{R}$, we extend it over $\mathbb{R}^{d}$ as
@@
 […]
+[…]
+We also define the normed spaces for $1\leq p\leq\infty$ as
 […]
@@
 converges weakly to $\mu$.
-First we see the proof of 5.3, assuming $5.$
+First, we see the proof of 5.3, assuming $5.$
 Proof of Theorem 5.3: Let $\{\omega_{n}\}_{n\geq 1}$ be the sequence satisfying 5.6. Consider $R_{n}=R^{\omega_{n}}_{n}$ as the following map between normed spaces,
@@
 […]
-By log-convexity of $l_{p}$ norms and Holder’s Inequality,
+By log-convexity of $l_{p}$ norms and Hölder’s Inequality,
 […]
@@
 6. Proof of Lemma 5.5:
-We assume the Uniqueness Principle[Theorem 5.1] to be true. We first deduce Lemma 5.5 under this assumption and then prove Theorem 5.1 in the next section
+We assume the Uniqueness Principle[Theorem 5.1] to be true. We first deduce Lemma 5.5 under this assumption and then prove Theorem 5.1 in the next section.
 Proof of Lemma 5.5 : Let $\omega\in\mathcal{S}_{n}$ and let $f:D_{n}\to[0,\infty)$. Let
@@
 […]
-To see the equivalence one notes the fact that $z(x)=Qf$.
-Proof of Lemma 6.2: To prove the Lemma, we consider the Monge-Ampere Operator for a given concave function $z$ given by
+To see the equivalence, one notes the fact that $z(x)=Qf$.
+Proof of Lemma 6.2: To prove the Lemma, we consider the Monge-Ampère Operator for a given concave function $z$ given by
 $Mz(x)=\prod_{i=1}^{d}\Delta_{i}z(x)$
@@
 For the other summand, observe that
-$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg].\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\penalty 10000\ ||R_{n}g||_{\infty}.$
-The last line follows from the Strong Markov property as
+$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}$
+The last line follows from the Strong Markov property since
 $\{\tau_{Kn}\leq n^{2}\}\in\mathcal{F}_{\tau_{Kn}}=\{A:\forall\,t\in\mathbb{N},\,\{\tau_{Kn}\leq t\}\cap A\in\mathcal{F}_{t+1}\}$
-$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right)$
+$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right).$
 Since $|X_{\mathscr{E}(\omega)}(n)|_{1}^{2}$ is a submartingale, we have by Doob’s inequality that
 $\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]=\mathbb{P}_{x}[\max_{0\leq m\leq n^{2}}|X_{{\mathscr{E}(\omega)}}(m)|_{1}^{2}\geq K^{2}n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}.$
-Further more observe that since $x\in T_{n}$ we must have $|x|^{2}_{1}\leq n^{2}d^{2}$ for every $x\in T_{n}$. Therefore, we have
+Furthermore, observe that since $x\in T_{n}$, we must have $|x|^{2}_{1}\leq n^{2}d^{2}$ for every $x\in T_{n}$. Therefore, we have
 $\sup_{x\in T_{n}}\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}\leq\frac{d^{2}+1}{K^{2}}.$
@@
 […]
-By construction, the assumptions of Lemma 5.5 are satisfied, thereby the proof.
-Therefore, it is now enough to prove Theorem 5.1. We shall in the next section also see how this leads to a proof of Theorem 3.5.
+By construction, the assumptions of Lemma 5.5 are satisfied, thereby completing the proof.
+Therefore, it is now enough to prove Theorem 5.1. We shall in the next section also see how this leads to a natural proof of Theorem 3.5.
 7. Proof of Theorem 5.1 and Theorem 3.5
-Proof of Theorem 5.1: We begin by noting that the class $\mathscr{A}(\mathscr{E},f)$ is non-empty. Indeed consider
+Proof of Theorem 5.1: We begin by noting that the class $\mathscr{A}(\mathscr{E},f)$ is non-empty. Indeed, consider
 $u(x)=n(n+1)-|x|_{1}(|x|_{1}+1).$
@@
 $\sum_{i=1}^{d}\mathscr{E}(\omega)(x_{0},e_{i})[z(x_{0}+e_{i})+z(x_{0}-e_{i})-2\beta]<-f(x_{0}).$
-It is again easy to show that
+It is again easy to show that.
 $z^{\prime}(x)=\begin{cases}\beta,&\text{if }x=x_{0},\\ z(x),&\text{otherwise}.\end{cases}$
@@
 $\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\leq\frac{c(\mathscr{E},x)^{d}}{2d}\sum_{i=1}^{d}\Delta_{i}u(x)$
-Denote by the $i$ the environment function which gives $\frac{1}{2d}$ on all the coordinates except the first one where it assigns 0.
-Therfore we have that
+Denote by $i$ the environment function which gives $\frac{1}{2d}$ on all the coordinates except the first one, where it assigns 0.
+Therefore, we have that
 $\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)\subset\mathscr{A}(\mathscr{E},f)$
-Denote by $\{X_{j}\}_{j\geq 0}$ to be the simple symmetric random walk. Then we have by Theorem 5.1 that
+Denote by $\{X_{j}\}_{j\geq 0}$ the simple symmetric random walk. Then we have by Theorem 5.1 that
 $\inf_{u\in\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)}u(x)\geq\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x)$
@@
 $\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}]\geq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))].$
-Similarly we note that
+Similarly, we note that
 $\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\geq\sum_{i=1}^{d}\Delta_{i}u(x).$
@@
 As above, one can show that
-$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}f(X_{j})]$
+$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]$
 Therefore, we have established the following result.
@@
 8. Preparation for the proof of Theorem 3.7:
-We begin this section by discussing the theory of Action of Linear Transformations. As mentioned earlier, for this proof we shall assume $\mathscr{E}=\operatorname{Id}$.
+We begin this section by discussing the theory of Action of Linear Transformations.
+As mentioned earlier, for this proof and from now on we shall assume $\mathscr{E}=\operatorname{Id}$ unless otherwise indicated.
 8.1. Action of Linear Transformations:
@@
 It is straightforward that such a map is invertible.
-The induced random walk under $T$ in $\mathbb{Z}^{d}$ started from $x\in\mathbb{Z}^{d}$ is given by
+The induced random walk under the transformation $T$ in $\mathbb{Z}^{d}$ started from $x\in\mathbb{Z}^{d}$ is given by
 $\mathbb{P}_{0}^{\omega}(T(X_{n})=T(X_{n-1})+T(e_{i}))=\mathbb{P}_{0}^{\omega}(X_{n}=X_{n-1}+e_{i})=\omega(X_{n-1},e_{i})$
-Call the transition kernel of the environment under this new random walk, i.e., the process $\tau_{T(X_{n})}\omega$, $K^{T}_{d}(\omega,B)$. Then we have the result,
+Call the transition kernel of the environment under this new random walk, i.e., the process $\tau_{T(X_{\omega}(n))}\omega$, $K^{T}_{d}(\omega,B)$. Then we have the result,
 Theorem 8.2.
@@
 Note that $T.\omega(o,e)=\omega(o,e)$ and by Theorem 8.1,
-$\{\tau_{e}{T.\omega}\in T(B)\}=\{T(\tau_{Tx}\omega)\in T(B)\}=\{\tau_{Tx}(B)\in B\}\quad\penalty 10000\ \forall e\in V_{d}.$
+$\{\tau_{e}{T.\omega}\in T(B)\}=\{T(\tau_{T.e}\omega)\in T(B)\}=\{\tau_{T.e}\omega\in B\}\quad\penalty 10000\ \forall e\in V_{d}.$
 This completes the proof. ∎
@@
 Corollary 8.3.
-If $\tau_{X_{\mathscr{E}(\omega)}(n)}\omega$ has a stationary measure $Q$, $\tau_{T(X_{\mathscr{E}(\omega)}(n))}\omega$ is also stationary under $Q\circ T$.
+If $\tau_{X_{\omega}(n)}\omega$ has a stationary ergodic measure $Q$, $\tau_{T(X_{\omega}(n))}\omega$ is also stationary and ergodic under $Q\circ T$.
 Proof.
@@
 $\mathbb{E}_{\mathbb{Q}}[\Phi]=\mathbb{Q}[h\in[a,b]]>0.$
-But this contradicts 3 since $\Phi=0$ ($\mathbb{Q}$ a.e.) Now we note that from the martingale property and the Markov property,
+But this contradicts 3 since $\Phi=0$ ($\mathbb{Q}$ a.e.). Now we note that from the martingale property and the Markov property and the fact that $\{h(\bar{\omega}(n)\}_{n\geq 0}$ is a martingale,
 $\boldsymbol{1}_{B}(\omega)=\mathbb{E}_{\mathbb{Q}}\Big[\,\boldsymbol{1}_{B}(\bar{\omega}(1))\,\big|\,\omega\,\Big]=K_{d}\,\boldsymbol{1}_{B}(\omega)\quad\mathbb{Q}\penalty 10000\ \text{-a.e.}$ (4)
@@
 The first part of the proof is to construct $\mathbb{Q}$.
-Let $\lambda^{\prime}$ be an invariant measure of $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 1}$. The first make an observation on this measure.
-Let $\alpha:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ given by $\alpha(e_{i})=-e_{i}\quad i=1,2,...,d$.
+Let $\lambda^{\prime}$ be an invariant measure of $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 0}$. First, we make an observation on this measure.
+Let $\alpha:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ given by $\alpha(e_{i})=-e_{i}\quad i=1,2,\dots,d$.
 Lemma 9.1.
@@
 Furthermore $\lambda^{\prime}\circ\alpha$ is mutually absolutely continuous to $\mu\circ\alpha$.
-One can also show that $\mu\circ\alpha$ is stationary ergodic with respect to the group-of-shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$.
+One can also show that $\mu\circ\alpha$ is stationary and ergodic with respect to the group-of-shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$.
 Let $g:\mathcal{S}(d)\to\mathbb{R}$ be a measurable function. By Birkhoff’s Ergodic Theorem,
@@
 $G_{d}(\omega,B):=K^{\alpha}_{d}(\omega,B)=K_{d}(\alpha.\omega,\alpha.B)\quad\forall B\in\mathcal{B}(d)^{\prime}.$
-We redefine $X_{\Lambda(\omega)}(n)$ in the following manner.
+We redefine $\{X_{\Lambda(\omega)}(n)\}_{n\geq 0}$ in the following manner.
 $L_{\Lambda(\omega)}g(X_{\Lambda(\omega)}(n))=\mathbb{E}[\gamma_{n}L_{\omega}g(X_{\Lambda(\omega)}(n))+(1-\gamma_{1})L_{\alpha(\omega)}g(X_{\Lambda(\omega)}(n))]\quad\forall g:\mathbb{Z}^{d}\to\mathbb{R},\penalty 10000\ \gamma_{n}\sim\text{Ber}\left(\frac{1}{2}\right)$
-where $L_{\Lambda(\omega)},L_{\omega},L_{\alpha(\omega)}$ are the generators of $X_{\Lambda(\omega)},X_{\omega},X_{\alpha(\omega)}$ respectively and $\gamma_{n}$ is a Bernoulli Random Variable independent of the environment and position of the walk.
+where $L_{\Lambda(\omega)},L_{\omega},L_{\alpha(\omega)}$ kernels of $X_{\Lambda(\omega)},X_{\omega},X_{\alpha(\omega)}$ respectively and $\gamma_{n}$ is a Bernoulli Random Variable independent of the environment and position of the walk.
 Further $\{\gamma_{n}\}_{n\geq 1}$ is a sequence of i.i.d.
@@
 We therefore define the conditional environment process
-$K_{d}^{\Lambda}(\omega,B)\Big|\gamma=K_{\gamma}(\omega,B)=\gamma K_{d}(\omega,B)+(1-\gamma)G_{d}(\omega,B).$
+$K_{d}^{\Lambda}(\omega,B)\Big|\gamma:=K_{\gamma}(\omega,B)=\gamma K_{d}(\omega,B)+(1-\gamma)G_{d}(\omega,B).$
 We now state an easy lemma without proof.
@@
 Let $n\geq 1$.
-Let $B_{0},B_{1},...,B_{n}\in\mathcal{B}(d)^{\prime}$.
-Let the probability measure associated to $\omega$ be $\lambda^{\prime}$.
+Let $B_{0},B_{1},\dots,B_{n}\in\mathcal{B}(d)^{\prime}$.
+Let the probability measure associated with $\omega$ be $\lambda^{\prime}$.
 Let $\bar{\omega}(n)=\tau_{X_{\Lambda(\omega)}(n)}\omega$.
@@
 To see how this implies stationarity, take $B_{i}=\pi_{0}^{-1}(A_{i})$ where $A_{i}\in\mathcal{B}(d)$. Then
-$\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{0}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})\Big|\gamma_{1},...,\gamma_{n}]=\frac{1}{2^{n}}\sum_{(\gamma_{1},...,\gamma_{n})\in\{0,1\}^{n}}\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{0}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})\Big|\gamma_{1},...,\gamma_{n}]$
-The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{1}),...\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$.
-Now conclude stationarity using the stationarity of $\bar{\omega}(n)$ and taking $\gamma_{i}=1$.
+$\lambda^{\prime}\left[\bigcap_{k=0}^{n}\{\bar{\omega}(k)\in\pi_{0}^{-1}(A_{k})\}\,\middle|\,\gamma_{1},\dots,\gamma_{n}\right]=\frac{1}{2^{n}}\sum_{\{\gamma_{i}\}_{i=1}^{n}\in\{0,1\}^{n}}\lambda^{\prime}\left[\bigcap_{k=0}^{n}\{\bar{\omega}(k)\in\pi_{0}^{-1}(A_{k})\}\,\middle|\,\{\gamma_{i}\}_{1\leq i\leq n}\right]$
+The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{1}),\dots,\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$.
+Now, conclude stationarity using the stationarity of $\bar{\omega}(n)$ and taking $\gamma_{i}=1$.
 It is clear that
-$\tau_{X_{\Lambda(\omega)}(n)}\omega|\{\gamma_{i}=1\quad i=1,2,....,n\}=\tau_{X_{\omega}(n)}\omega.$
-Therefore it is now enough to establish Lemma 9.3.
+$\tau_{X_{\Lambda(\omega)}(n)}\omega|\{\gamma_{i}=1\quad i=1,2,\dots,n\}=\tau_{X_{\omega}(n)}\omega.$
+Therefore, it is now enough to establish Lemma 9.3.
 Proof of Lemma 9.3: We shall prove this result using induction. First for $n=1$, we note that
@@
 $\int_{B_{0}}K_{1-\gamma}(\omega,\alpha.B_{1})d\lambda^{\prime}\circ\alpha=\lambda^{\prime}(\alpha(\omega)\in B_{0},\bar{\omega}(1)\in\alpha.B_{1}|1-\gamma)$
-Suppose the result is true for $n=m$. Let $n=m+1$. Then we must verify the statement for $k=m+1$ only. For $k=m+1$, we write
-$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times\alpha.B_{m}}K_{\gamma_{m+1}}(\omega_{m},B_{m+1})d\lambda^{\prime}(\omega_{m},...,\omega|\gamma_{1},...,\gamma_{m})$
-Perform the substitution $(\omega_{m},...,\omega)\mapsto(\alpha.\omega_{m},...,\omega)$ to get
-$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times B_{m}}K_{\gamma_{m+1}}(\alpha.\omega_{m},B_{m+1})d\lambda^{\prime}(\alpha.\omega_{m},...,\omega|\gamma_{1},...,\gamma_{m})$
+Suppose the result is true for $n=m$. Let $n=m+1$. We verify the statement for $k=m+1$ only for simplicity as others follow from the technique of proof. For $k=m+1$, we write
+$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times\alpha.B_{m}}K_{\gamma_{m+1}}(\omega_{m},B_{m+1})d\lambda^{\prime}(\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m})$
+Perform the substitution $(\omega_{m},\dots,\omega)\mapsto(\alpha.\omega_{m},\dots,\omega)$ to get
+$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times B_{m}}K_{\gamma_{m+1}}(\alpha.\omega_{m},B_{m+1})d\lambda^{\prime}(\alpha.\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m})$
 Using Lemma 9.2, the integral reduces to
@@
 Now we write the above expression as
-$\mathbb{E}\Big[\mathbb{E}\Big[\boldsymbol{1}_{\alpha.B_{m+1}}(\bar{\omega}(m+1))\Big|\bar{\omega}(m),1-\gamma_{m+1}\Big]\boldsymbol{1}_{\bar{\omega}(m)}(\alpha.B_{m})\boldsymbol{1}_{\bar{\omega}(m-1)}(B_{m-1})\dots,\boldsymbol{1}_{\omega}(B_{0})\Big|\gamma_{1},...,\gamma_{m}\Big]$
+$\mathbb{E}\Big[\mathbb{E}\Big[\boldsymbol{1}_{\alpha.B_{m+1}}(\bar{\omega}(m+1))\Big|\bar{\omega}(m),1-\gamma_{m+1}\Big]\boldsymbol{1}_{\bar{\omega}(m)}(\alpha.B_{m})\boldsymbol{1}_{\bar{\omega}(m-1)}(B_{m-1})\dots,\boldsymbol{1}_{\omega}(B_{0})\Big|\gamma_{1},\dots,\gamma_{m}\Big]$
 This reduces by definition to the desired form, which completes the proof.
-Acknowledgements:
-The author is deeply indebted to Dr.
-Soumendu Sundar Mukherjee of Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for helpful discussions and his immense help with the manuscript.
-The author also thanks his classmates Himadri Mandal and Aryama Ghosh for their helpful comments on the exposition and the manuscript in general.

### v3

@@
 Roughly speaking, it says there is a deterministic limiting velocity of the walk.
-Thid have been studied in detail in $d=1$ by Solomon[11] in 1975.
+This has been studied in detail in $d=1$ by Solomon[10] in 1975.
 In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
@@
 Does there exist a deterministic vector $v$ such that
-$\frac{X_{n}}{n}\to v\ \ \ \ \text{a.e.}$
+$\frac{X_{n}}{n}\to v\quad\text{a.e.}$
 for almost every choice of environment?
@@
 Conjecture 1.2.
-Does every balanced, elliptic(probability of going along every axis at every point is positive) walk in $\mathbb{Z}^{d}$ recurrent in $d=1,2$ and transient in $d\geq 3$?
+Is every balanced, elliptic(probability of going along every axis at every point is positive) walk in $\mathbb{Z}^{d}$ recurrent in $d=1,2$ and transient in $d\geq 3$?
 We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest.
@@
 It is important to note that this section shall not deal with the notion of a Random Environment.
-This is one of the key contributions of this p, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
+This is one of the key contributions of this paper, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
 Our other main contribution will be to deduce the ergodicity of a certain stochastic process, which we will call the local process.
@@
 To the best of our knowledge, these theorems are novel.
-Our results hold for all uniformly elliptic RWRE, and the conditions we impose study the extent to which the condition of uniform ellipticity can be dropped.
+Our results hold for all uniformly elliptic RWRE in dimensions equal to or greater than one, and the conditions we impose study the extent to which the condition of uniform ellipticity can be dropped.
 Acknowledgements:
-The author is an undergraduate student at the Indian Statistical Insitute, Kolkata.
-This work was completed during the third-year of his studies.
+The author is an undergraduate student at the Indian Statistical Institute, Kolkata.
+This work was completed during the third year of his studies.
 He is deeply indebted to Dr.
-Soumendu Sundar Mukherjee of Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for immensely helpful discussions and his substantial help with the manuscript.
+Soumendu Sundar Mukherjee of the Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Kolkata, for immensely helpful discussions and his substantial help with the manuscript.
 The author also thanks his classmates Himadri Mandal and Aryama Ghosh for their helpful comments on the exposition and the manuscript in general.
@@
 Before stating the main results, we formally define the setting we work under. We adopt the formalism as in [7].
-Define by $\mathbb{Z}^{d}$ the integer lattice and $\{e_{i}\}_{i=1,..,d}$ the unit vectors. Define the set
-$V_{d}=\{\pm e_{i}:i=1,\dots,d\}\cup\{0\}$
+Define by $\mathbb{Z}^{d}$ the integer lattice and $\{e_{i}\}_{i=1,\dots,d}$ the unit vectors. Define the set
+$V_{d}=\{\pm e_{i}:i=1,\dots,d\}\cup\{0\}.$
 Define the following sets:
@@
 Let $\mathcal{B}(d)$, $\mathcal{B}(d)^{\prime}$ be the Borel-$\sigma$ algebra on $S(d)$, $\mathcal{S}(d)$ respectively.
-We call any measurable $\mathscr{E}:\mathcal{S}(d)\to\mathcal{S}(d)$ as the environment function.
+Take $\mathscr{E}:\mathcal{S}(d)\to\mathcal{S}(d)$ such that $\mathscr{E}(\omega)(x)\equiv\mathscr{E}(\omega(x)):S(d)\to S(d)$ is a measurable function on $(S(d),\mathcal{B}(d))$.
+We call this the environment function.
 This function essentially contains the information on the transition probabilities of the random walk based on the function $\omega$.
@@
 Let $\mu$ be a probability measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. We call the environment function balanced if
-$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega(x))=\mathscr{E}_{i+d}(\omega(x))\quad i=1,2,\dots,d.$
+$\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega)(x)=\mathscr{E}_{i+d}(\omega)(x)\quad i=1,2,\dots,d.$
 We shall call the environment elliptic if
-$\mu[\mathscr{E}_{k}(\omega)(x)>0:\forall x\in\mathbb{Z}^{d},\forall k]=1$
+$\mu[\mathscr{E}_{k}(\omega)(x)>0:\forall x\in\mathbb{Z}^{d},\forall k]=1.$
 Define the group of shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$ on $\omega\in\mathcal{S}(d)$ given by
@@
 For $\mathscr{E}(\omega)\in\mathcal{S}(d)$, $v\in V_{d}$, $\mathscr{E}$ an environment function
-$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}_{i}(\omega)(x)\quad\text{if}\penalty 10000\ v=e_{i}$
-$\mathscr{E}(\omega)(x,-v)$ $=\mathscr{E}_{i+d}(\omega)(x)\quad\text{if}\penalty 10000\ v=-e_{i}$
-$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}_{0}(\omega)\quad\text{if}\penalty 10000\ v=0.$
+$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}(x,v)=\mathscr{E}_{i}(\omega)(x)\quad\text{if}\penalty 10000\ v=e_{i}$
+$\mathscr{E}(\omega)(x,-v)$ $=\mathscr{E}(x,v)=\mathscr{E}_{i+d}(\omega)(x)\quad\text{if}\penalty 10000\ v=-e_{i}$
+$\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}(x,v)=\mathscr{E}_{0}(\omega)\quad\text{if}\penalty 10000\ v=0.$
 We shall assume $\mu$ to be stationary and ergodic with respect to the group of shifts $\tau_{x}$. More precisely,
-$\mu[\tau_{-x}\omega\in A]=\mu[\omega\in A]\quad\forall A\in\mathcal{B}(d)^{\prime},\forall x\in\mathbb{Z}^{d}$
-$\forall B\in\mathcal{B}(d)^{\prime},\quad\mu\left[\cap_{x\in\mathbb{Z}^{d}}\tau_{x}B\right]\in\{0,1\}$
+$\mu[\tau_{-x}\omega\in A]=\mu[\omega\in A]\quad\forall A\in\mathcal{B}(d)^{\prime},\forall x\in\mathbb{Z}^{d},$
+$\forall B\in\mathcal{B}(d)^{\prime},\quad\mu\left[\cap_{x\in\mathbb{Z}^{d}}\tau_{x}B\right]\in\{0,1\}.$
 As a special case, $\mu$ is said to be an independent and identically distributed (i.i.d.) law if $\{\omega(x)\}_{x\in\mathbb{Z}^{d}}$ is an i.i.d.
@@
 Define the Markov Chain of the Random Walk $\{X_{\mathscr{E}(\omega)}(k)\}_{k\geq 0}$ in $\mathbb{Z}^{d}$ by the transition probabilities,
-$\mathbb{P}[X_{\mathscr{E}(\omega)}(k+1)$ $=x+v|X_{\mathscr{E}(\omega)}(k)=x]=\mathscr{E}(x,v)\quad\forall v\in V_{d}$
-Therefore, the kernel of this Markov Chain are given by
+$\mathbb{P}[X_{\mathscr{E}(\omega)}(k+1)$ $=x+v|X_{\mathscr{E}(\omega)}(k)=x]=\mathscr{E}(x,v)\quad\forall v\in V_{d}.$
+Therefore, the kernel of this Markov Chain is given by
 $L_{\mathscr{E}(\omega)}(g(x))=\sum_{v\in V_{d}}\mathscr{E}(\omega)(x,v)g(x+v),$
@@
 Assumption 2.2.
-Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d))$. Define the quantity $c(\mathscr{E},x)$ as follows
-$c(\mathscr{E},x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}$
+Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. Define the quantity $c(\mathscr{E},x)$ as follows
+$c(\mathscr{E},x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}.$
 Then we assume that
@@
 We use the notion of concave functions defined on the compact $L_{1}$ ball $B_{1}[0,r]\cap\mathbb{Z}^{d}$. A function $z:B_{1}[0,r]\cap\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ $(r>0)$ is said to be concave if
-$z(x+v)+z(x-v)-2z(x)<0\quad x\in\operatorname{int}B_{1}[0,r]\cap\mathbb{Z}^{d}\quad\forall v\in V_{d}\setminus\{0\}$
+$z(x+v)+z(x-v)-2z(x)\leq 0\quad x\in\operatorname{int}B_{1}[0,r]\cap\mathbb{Z}^{d}\quad\forall v\in V_{d}\setminus\{0\},$
 where $\operatorname{int}$ denotes the topological interior of a set.
@@
 Theorem 3.1.
-Let $d\geq 2$. Assume $\mu$ is a stationary, ergodic Law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a balanced environment function satisfying assumptions 2.2 and 2.3.
+Let $d\geq 1$. Assume $\mu$ is a stationary, ergodic Law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a measurable balanced environment function satisfying assumptions 2.2 and 2.3.
 Then there exists a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ such that the environment process $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ is stationary, ergodic under this measure and is mutually absolutely continuous to $\mu$.
@@
 Corollary 3.2.
+Let $d\geq 1$.
 Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ and $\mathscr{E}$ a measurable balanced environment function.
@@
 $\frac{X_{\mathscr{E}(\omega)}([nt])}{\sqrt{n}}$
-converges weakly to the standard Brownian Motion on $\mathbb{R}^{d}$ with non degenerate covariance matrix of the form $(b_{i}\delta_{ij})_{1\leq i,j\leq d}$ for $\omega$-$\mu$ a.e.
+converges weakly to the standard Brownian Motion on $\mathbb{R}^{d}$ with non degenerate covariance matrix of the form $(b_{i}\delta_{ij})_{1\leq i,j\leq d}$ for $\omega$ $\mu$-a.e.
 These results notably weaken the assumptions of ellipticity and derive an alternate proof of Theorem 1 of [4] for $d\geq 1$, which states that
@@
 Then the Green’s function $G^{\mathscr{E}(\omega)}:\mathbb{Z}^{d}\times\mathbb{Z}^{d}\to\mathbb{R}^{+}\cup\{0\}$ is given by
-$\mathbb{E}_{x}[\sum_{j=0}^{\infty}\boldsymbol{1}_{X_{\mathscr{E}(\omega)}=y}]$
+$G^{\mathscr{E}(\omega)}(x,y)=\mathbb{E}_{x}[\sum_{j=0}^{\infty}\boldsymbol{1}_{X_{\mathscr{E}(\omega)}(j)=y}],$
 where the expectation is taken conditional on $X_{\mathscr{E}(\omega)}(0)=x$. For the simple symmetric random walk, we shall denote its Green’s function as just $G$.
@@
 Let $d\geq 1$. Let $(\omega,\mathscr{E})$ be any balanced elliptic environment. Then
-$\frac{1}{2d}G(x,x)\leq G^{\mathscr{E}(\omega)}(x,x)\leq\frac{1}{c(\mathscr{E},x)^{d}}G(x,x)$
+$\frac{1}{2d}G(x,x)\leq G^{\mathscr{E}(\omega)}(x,x)\leq\frac{1}{c(\mathscr{E},x)^{d}}G(x,x).$
 This shows the conclusions of Pölya’s Theorem, which states
@@
 Theorem 3.8.
-Let $d\geq 1$ and let $\mu$ a stationary ergodic environment law on $\mathcal{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
+Let $d\geq 1$, and let $\mu$ a stationary ergodic environment law on $\mathcal{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
   1. (i)
@@
     $\frac{X_{\omega}([nt])-\sum_{i=0}^{[nt]}d(0,\tau_{X_{\omega}(m)}\omega)}{\sqrt{n}}$
-    converges to the standard Brownian Motion on $\mathbb{R}^{d}$ with non-degenerate covariance matrix $\omega$-$\mu$ a.e.
+    converges to the standard Brownian Motion on $\mathbb{R}^{d}$ with non-degenerate covariance matrix $\omega$ $\mu$-a.e.
 3.2. Remarks:
@@
 This allows us to “couple” the original walk and a reflected walk.
-This technique allows us to control the paths of the random walk which shall be key to the proof.
+This technique allows us to control the paths of the random walk, which shall be key to the proof.
 The control argument will therefore allow us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process.
@@
   1. 1.
-    $u\equiv 0$ on $\partial D_{n}$.
+    $u\equiv 0$ on $\partial D_{n}$,
   2. 2.
-    $L_{\mathscr{E}(\omega)}u(x)\leq-f(x)\quad\forall x\in D_{n}^{\circ}$.
+    $L_{\mathscr{E}(\omega)}u(x)\leq-f(x)+u(x)\quad\forall x\in D_{n}^{\circ}$.
 Theorem 5.1 (Uniqueness Principle).
 Let $(\omega,\mathscr{E})$ be any balanced, elliptic environment.
-There exists an unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that $L_{\mathscr{E}(\omega)}z(x)=-f(x)\quad\forall x\in D_{n}^{\circ}$ given by
-$z(x)=\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)})(x)]$
+There exists an unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that $L_{\mathscr{E}(\omega)}z(x)=-f(x)+z(x)\quad\forall x\in D_{n}^{\circ}$ given by
+$z(x)=\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)})(x)],$
 where the expectation is taken assuming $X_{\mathscr{E}(\omega)}(0)=x\in D_{n}^{\circ}$. Furthermore, this function is also given alternately by
-$z(x)=\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x),\quad\forall x\in\mathbb{Z}^{d}.$
+$z(x)=\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x),\quad\forall x\in D_{n}.$
 Now we return to Theorem 3.1 and Corollary 3.2. Assume without loss of generality, $X_{\mathscr{E}(\omega)}(0)=0$.
@@
 Define $\tau_{X_{\mathscr{E}(\omega)}(j)}\omega=Y_{j}$. Then
-$\mathbb{P}[Z_{j}^{i}=e_{i}]=\mathbb{P}_{x}[Z_{j}^{i}=-e_{i}]=\mathscr{E}_{i}(Y_{j}(0))$
+$\mathbb{P}[Z_{j}^{i}=e_{i}]=\mathbb{P}_{x}[Z_{j}^{i}=-e_{i}]=\mathscr{E}_{i}(Y_{j})(0).$
 Under the filtration
@@
 $\lim_{n\to\infty}\frac{1}{n}\sum_{k=0}^{n-1}\mathscr{E}[\omega(X_{\mathscr{E}(\omega)}(k))]=b.$
-for $\omega$-$\mu\penalty 10000\ \text{a.e.}$
+for $\omega$ $\mu\text{-a.e.}$
 We prove this by finding an appropriate measure mutually absolutely continuous to $\mu$ such that the ergodic theorem holds.
 Recall that $Y_{j}$ is a Markov chain with the kernel
-$K^{\mathscr{E}}_{d}(g(Y_{j}))=\sum_{i=1}^{d}\mathscr{E}_{i}(Y_{j})(0)[g(\tau_{e_{i}}Y_{j})+g(\tau_{-e_{i}}Y_{j})]$
+$K^{\mathscr{E}}_{d}(g(Y_{j}))=\sum_{i=1}^{d}\mathscr{E}_{i}(Y_{j})(0)[g(\tau_{e_{i}}Y_{j})+g(\tau_{-e_{i}}Y_{j})]+\mathscr{E}_{0}(Y_{j})(0)g(Y_{j})$
 for all $g:\mathcal{S}(d)\to\mathbb{R}^{d}$. Therefore 3.2 is equivalent to the following theorem,
@@
 Definition 5.4 (Periodic Environment).
-We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$.
-[…]
-$\hat{\omega}(x)$ $=\omega([x]_{\sim}),$
-$\mathscr{E}(\hat{\omega})(x)$ $=\mathscr{E}(\omega)([x]_{\sim})$
-where $x\in\mathbb{Z}^{d}$. Let $\mathcal{S}_{n}$ denote the set of all such $\hat{\omega}$. Note that $\mathcal{S}_{n}\subset\mathcal{S}(d)$.
-The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the kernel $L_{\mathscr{E}(\hat{\omega})}$.
+We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$. Let $\mathcal{S}_{n}$ denote the set of all such $\omega$.
+Fix $\omega\in\mathcal{S}(d)$. We take the periodic environment $(\omega_{n},\mathscr{E})$ as
+$\omega_{n}(x)$ $=\omega([x]_{\sim})$
+$\mathscr{E}(\omega_{n})(x)$ $=\mathscr{E}(\omega)([x]_{\sim}),$
+where $x\in\mathbb{Z}^{d}$.
+The corresponding random walk $\{\hat{X}_{n}\}$ in this periodic environment has the kernel $L_{\mathscr{E}(\omega_{n})}$.
 For a function $g:T_{n}\to\mathbb{R}$, we extend it over $\mathbb{R}^{d}$ as
@@
 $R_{n}^{\omega}g(x)=\sum_{j=0}^{\infty}(1-\frac{1}{n^{2}})^{j}L_{\mathscr{E}({\omega})}^{j}g(x)$
-If $g:T_{n}\to\mathbb{R}$. Define the $l_{p}$($1\leq p\leq\infty)$ norms with respect to the normalized counting measure on $T_{n}$ by
+For $g:T_{n}\to\mathbb{R}$, define the $l_{p}$($1\leq p\leq\infty)$ norms with respect to the normalized counting measure on $T_{n}$ by
 […]
@@
 […]
-To show that 5.5 $\implies$ 5.3, we first state a lemma by Parthsarathy[9] under the assumption that $\mu$ is stationary.
+To show that 5.5 $\implies$ 5.3, we first state the so called multidimensional ergodic theorem[Chapter 6, Theorem 2.8 of [5]]
 Lemma 5.6.
-[Lemma $5$ of [7]] For each $n$ there exists $\omega_{n}\in\mathcal{S}_{n}$ such that the sequence of measures $\mu_{n}$ given by
-$\mu_{n}=\frac{1}{(2n)^{d}}\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}$
-converges weakly to $\mu$.
+For each $n\geq 1$, define the sequence of measures $\mu_{n}$,
+$\mu_{n}=\frac{1}{(2n)^{d}}\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}.$
+Then we have:
+  1. 1.
+    $\mu_{n}\to\mu$ weakly $\mu$-a.e.
+  2. 2.
+    For every measurable $f:S(d)\to\mathbb{R}$ such that $f\in L^{p}(\mu)$ where $p\geq 1$, $\mu_{n}f\to\mu f$ $\mu$-a.e.
 First, we see the proof of 5.3, assuming $5.$
-Proof of Theorem 5.3: Let $\{\omega_{n}\}_{n\geq 1}$ be the sequence satisfying 5.6. Consider $R_{n}=R^{\omega_{n}}_{n}$ as the following map between normed spaces,
+Proof of Theorem 5.3: Consider $R_{n}=R^{\omega_{n}}_{n}$ as the following map between normed spaces,
 $R_{n}:l^{d}(T_{n})\to l^{\infty}(T_{n}).$
 Then the dual of $R_{n}$ is the map
-$R^{*}_{n}:l^{1}(T_{n})\to l^{\frac{d}{d-1}}(T_{n})$
-Let the invariant measure for the random walk on the group $T_{n}$ given by $X_{\mathscr{E}(\omega_{n})}(k)]_{\sim}$ be $\frac{\phi_{n}(x)}{(2n)^{d}}$.
+$R^{*}_{n}:l^{1}(T_{n})\to l^{\frac{d}{d-1}}(T_{n}).$
+Let the invariant measure for the random walk on the group $T_{n}$ given by $\{[X_{\mathscr{E}(\omega_{n})}(k)]_{\sim}\}_{k\geq 0}$ be $\frac{\phi_{n}(x)}{(2n)^{d}}$.
 Note that $\phi_{n}$ is the density with respect to the normalised counting measure on $T_{n}$.
 Therefore, one can check (or refer to Section 2 of [4]) the invariant measure for $\{\omega_{n}(\hat{X}_{k})\}_{k\geq 0}$ is given by
-$\lambda_{n}(x)=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
+$\lambda_{n}(\tau_{x}\omega_{n})=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
 Since $\mathcal{S}(d)$ is compact with respect to the product topology, there exists a subsequence $n_{k}$ such that
@@
 […]
-Since
+By Lemma 5.6,
 […]
 This shows
-$\frac{d\lambda_{n}}{d\mu_{n}}||_{\frac{p}{p-1}}<\infty\quad\mu\penalty 10000\ \text{-a.e.}$
-Since $\lambda_{n}$ extends as a measurable function over the entire $\mathbb{Z}^{d}$, we have that the family of functions
+$\frac{d\lambda_{n}}{d\mu_{n}}||_{\frac{p}{p-1}}<\infty\quad\mu\text{-a.e.}$
+Since $\lambda_{n}$ extends as a measurable function over the entire $\mathcal{S}(d)$, we have that the family of functions
 $\big\{\frac{d\lambda_{n}}{d\mu_{n}}\big\}_{n\geq 1}$
 is uniformly integrable with respect to $\mu$. Therefore by Dunford-Pettis Theorem, $\lambda\ll\mu$ and
-$\int_{\mathcal{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<\infty$
+$\int_{\mathcal{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<\infty.$
 Referring to arguments from [3], $\lambda$ is ergodic and $\lambda\ll\mu$ and $\lambda\gg\mu$. This concludes the proof.
@@
 $\tau_{n}$ $=\inf\{m:X_{\mathscr{E}(\omega)}(m)\in\partial D_{n}\}$
-$Qf(x)$ $=\mathbb{E}_{x}[\sum_{k=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(k))]$
+$Qf(x)$ $=\mathbb{E}_{x}[\sum_{k=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(k))],$
 where the expectation is taken assuming $X_{{\mathscr{E}(\omega)}}(0)=x\in D_{n}^{\circ}$
@@
 for some $c_{2}\equiv c_{2}(d)>1$.
-Lemma 6.2 is equivalent to the Lemma 6.1 which states that,
+Lemma 6.1 is equivalent to Lemma 6.2 which states that,
 Lemma 6.2.
@@
 […]
-To see the equivalence, one notes the fact that $z(x)=Qf$.
+To see the equivalence, one notes the fact that $z(x)=Qf(x)$.
 Proof of Lemma 6.2: To prove the Lemma, we consider the Monge-Ampère Operator for a given concave function $z$ given by
-$Mz(x)=\prod_{i=1}^{d}\Delta_{i}z(x)$
+$Mz(x)=\prod_{i=1}^{d}\Delta_{i}z(x).$
 Note that from Theorem 5.1 one has that
-$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)=f(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x)$
+$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)=f(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x).$
 Define the sets for $x\in D_{n}^{\circ}$
@@
 […]
-Choose $c_{3}=c_{4}\vee 1$. This completes the proof of 6.2(c).
+Choose $c_{3}=c_{4}\vee 2$. This completes the proof of 6.2(c).
 Now to bound the resolvent[recall the statement of Lemma 5.5], first we take $m=Kn$ such that $T_{n}\subset D_{m}$.
@@
 For the other summand, observe that
-$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}$
+$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}.$
 The last line follows from the Strong Markov property since
@@
 Furthermore, observe that since $x\in T_{n}$, we must have $|x|^{2}_{1}\leq n^{2}d^{2}$ for every $x\in T_{n}$. Therefore, we have
-$\sup_{x\in T_{n}}\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}\leq\frac{d^{2}+1}{K^{2}}.$
+$\sup_{x\in T_{n}}\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\leq\sup_{x\in T_{n}}\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}=\frac{d^{2}+1}{K^{2}}.$
 Now choose $K$ such that
-$\frac{d^{2}+1}{K^{2}}$ $<\frac{1}{2}-\frac{1}{e},\quad T_{n}\subset D_{m}$
+$\frac{d^{2}+1}{K^{2}}$ $<\frac{1}{2}-\frac{1}{e},\quad T_{n}\subset D_{m}.$
 Therefore, combining the deductions we have that
@@
 We claim that this function satisfies the second condition of the Theorem. Indeed if it does not then $\exists x_{0}\in D_{n}^{\circ}$ such that
-$L_{\mathscr{E}(\omega)}u(x_{0})<-f(x_{0}).$
-One can choose $0<\beta<f(x_{0})$ such that
+$L_{\mathscr{E}(\omega)}z(x_{0})<-f(x_{0})+z(x_{0}).$
+One can choose $0<\beta<z(x_{0})$ such that
 $\sum_{i=1}^{d}\mathscr{E}(\omega)(x_{0},e_{i})[z(x_{0}+e_{i})+z(x_{0}-e_{i})-2\beta]<-f(x_{0}).$
-It is again easy to show that.
+It is again easy to show that
 $z^{\prime}(x)=\begin{cases}\beta,&\text{if }x=x_{0},\\ z(x),&\text{otherwise}.\end{cases}$
@@
 One also notes that for every $j\geq 1$,
-$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]=\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\omega}(k))$
+$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]=\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\mathscr{E}(\omega)}(k)).$
 Letting $k\to\infty$ completes the proof.
@@
 Observe that for every concave function $u\in\mathscr{A}(\mathscr{E},f)$ one has
-$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\leq\frac{c(\mathscr{E},x)^{d}}{2d}\sum_{i=1}^{d}\Delta_{i}u(x)$
+$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\leq\frac{c(\mathscr{E},x)^{d}}{2d}\sum_{i=1}^{d}\Delta_{i}u(x).$
 Denote by $i$ the environment function which gives $\frac{1}{2d}$ on all the coordinates except the first one, where it assigns 0.
 Therefore, we have that
-$\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)\subset\mathscr{A}(\mathscr{E},f)$
+$\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)\subset\mathscr{A}(\mathscr{E},f).$
 Denote by $\{X_{j}\}_{j\geq 0}$ the simple symmetric random walk. Then we have by Theorem 5.1 that
-$\inf_{u\in\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)}u(x)\geq\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x)$
+$\inf_{u\in\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)}u(x)\geq\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x).$
 or by using the other form,
@@
 This shows that
-$\mathscr{A}(\mathscr{E},f)\subset\mathscr{A}\left(i,\frac{f(x)}{2d}\right)$
+$\mathscr{A}(\mathscr{E},f)\subset\mathscr{A}\left(i,\frac{f(x)}{2d}\right).$
 As above, one can show that
-$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]$
+$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})].$
 Therefore, we have established the following result.
@@
 Let $f:D_{n}\to[0,\infty)$ be a function. Let $(\omega,\mathscr{E})$ be a balanced elliptic environment. Then for all $x\in D_{n}^{\circ}$, we have the inequality,
-$\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}]$
+$\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}].$
 One can now conclude Theorem 3.5 by taking the function $f:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$
@@
 We begin this section by discussing the theory of Action of Linear Transformations.
-As mentioned earlier, for this proof and from now on we shall assume $\mathscr{E}=\operatorname{Id}$ unless otherwise indicated.
+As mentioned earlier, for this proof and from now on, we shall assume $\mathscr{E}=\operatorname{Id}$ unless otherwise indicated.
 8.1. Action of Linear Transformations:
@@
 Theorem 8.1.
-$\tau_{x}T.\omega=T(\tau_{Tx}\omega)$
+$\tau_{x}T.\omega=T(\tau_{Tx}\omega).$
 Proof.
@@
 The induced random walk under the transformation $T$ in $\mathbb{Z}^{d}$ started from $x\in\mathbb{Z}^{d}$ is given by
-$\mathbb{P}_{0}^{\omega}(T(X_{n})=T(X_{n-1})+T(e_{i}))=\mathbb{P}_{0}^{\omega}(X_{n}=X_{n-1}+e_{i})=\omega(X_{n-1},e_{i})$
+$\mathbb{P}_{0}^{\omega}(T(X_{n})=T(X_{n-1})+T(e_{i}))=\mathbb{P}_{0}^{\omega}(X_{n}=X_{n-1}+e_{i})=\omega(X_{n-1},e_{i}).$
 Call the transition kernel of the environment under this new random walk, i.e., the process $\tau_{T(X_{\omega}(n))}\omega$, $K^{T}_{d}(\omega,B)$. Then we have the result,
 Theorem 8.2.
-$K_{d}^{\operatorname{Id}}\equiv K_{d}(T.\omega,T(B))=K^{T}_{d}(\omega,B)$ for every $B\in{\mathcal{B}(d)}$
+$K_{d}^{\operatorname{Id}}\equiv K_{d}(T.\omega,T(B))=K^{T}_{d}(\omega,B)$ for every $B\in{\mathcal{B}(d)^{\prime}}$.
 Proof.
@@
 $h(\bar{\omega}(n))$ is a $\tilde{P}_{\mathbb{Q}}$ martingale since
-$\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}\circ T^{n}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=h(\bar{\omega}(n))\quad\mathbb{Q}\penalty 10000\ \text{-a.e.}$
+$\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}\circ T^{n}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=h(\bar{\omega}(n))\quad\mathbb{Q}\text{-a.e.}$
 By the Levy Upwards Theorem,
-$h(\bar{\omega}(n))\to\boldsymbol{1}_{A}\qquad(n\to\infty)\quad\mathbb{Q}\penalty 10000\ \text{-\text{a.e.}}$ (3)
+$h(\bar{\omega}(n))\to\boldsymbol{1}_{A}\qquad(n\to\infty)\quad\mathbb{Q}\text{-a.e.}$ (3)
 Next we show $h=\boldsymbol{1}_{B}$($\mathbb{Q}$ a.e.) for some $B\in\mathcal{B}(d)^{\prime}$. If not then $\mathbb{Q}(h\in[a,b])>0$ for some $0<a<b<1$.
 But by Birkhoff’s Ergodic Theorem,
-$\frac{1}{n}\sum_{k=0}^{n-1}\boldsymbol{1}\{h(\bar{\omega}(k))\in[a,b]\}\to\Phi=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}\{h(\omega)\in[a,b]\}|\mathcal{I}]\quad\mathbb{Q}\penalty 10000\ \text{-a.e.}$
+$\frac{1}{n}\sum_{k=0}^{n-1}\boldsymbol{1}\{h(\bar{\omega}(k))\in[a,b]\}\to\Phi=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}\{h(\omega)\in[a,b]\}|\mathcal{I}]\quad\mathbb{Q}\text{-a.e.}$
 where $\mathcal{I}$ is the $\sigma$-algebra of invariant events.
@@
 But this contradicts 3 since $\Phi=0$ ($\mathbb{Q}$ a.e.). Now we note that from the martingale property and the Markov property and the fact that $\{h(\bar{\omega}(n)\}_{n\geq 0}$ is a martingale,
-$\boldsymbol{1}_{B}(\omega)=\mathbb{E}_{\mathbb{Q}}\Big[\,\boldsymbol{1}_{B}(\bar{\omega}(1))\,\big|\,\omega\,\Big]=K_{d}\,\boldsymbol{1}_{B}(\omega)\quad\mathbb{Q}\penalty 10000\ \text{-a.e.}$ (4)
+$\boldsymbol{1}_{B}(\omega)=\mathbb{E}_{\mathbb{Q}}\Big[\,\boldsymbol{1}_{B}(\bar{\omega}(1))\,\big|\,\omega\,\Big]=K_{d}\,\boldsymbol{1}_{B}(\omega)\quad\mathbb{Q}\text{-a.e.}$ (4)
 Combining 4 with ellipticity,
-$\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{e}\quad(\mu\text{-\text{a.e.} or }\mathbb{Q}\penalty 10000\ \text{-\text{a.e.}})\quad\forall\,e\in{V}_{d}.$ (5)
+$\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{e}\quad(\mu\text{-\text{a.e.} or }\mathbb{Q}\text{-a.e.})\quad\forall\,e\in{V}_{d}.$ (5)
 Iterating 5, we obtain
-$\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{x}\quad(\mu\text{-\text{-a.e.} or }\mathbb{Q}\penalty 10000\ \text{-\text{a.e.}})\quad\forall\,x\in\mathbb{Z}^{d}.$
+$\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{x}\quad(\mu\text{\text{-a.e.} or }\mathbb{Q}\text{\text{-a.e.}})\quad\forall\,x\in\mathbb{Z}^{d}.$
 Since $\mu[\omega\in B]=\mu[\tau_{x}\omega\in B]$ for every $x\in\mathbb{Z}^{d}$,
-$\boldsymbol{1}_{B}=\boldsymbol{1}_{B}\circ\tau_{x}\quad\mu\penalty 10000\ \text{\text{-a.e.} or}\penalty 10000\ \mathbb{Q}\penalty 10000\ \text{-\text{a.e.}}$ (6)
+$\boldsymbol{1}_{B}=\boldsymbol{1}_{B}\circ\tau_{x}\quad\mu\text{\text{-a.e.} or}\penalty 10000\ \mathbb{Q}\text{\text{-a.e.}}$ (6)
 Consider the set
@@
 $\lambda^{\prime}$ is an $\alpha$-invariant measure. More precisely,
-$\lambda^{\prime}\circ\alpha(B)=\lambda^{\prime}(B)\quad\forall B\in\mathcal{B}(d)^{\prime}$
+$\lambda^{\prime}\circ\alpha(B)=\lambda^{\prime}(B)\quad\forall B\in\mathcal{B}(d)^{\prime}.$
 Proof.
@@
 Recall that $\{X_{\omega}(n)\}_{n\geq 0}$ is the random walk with the law
-$\mathbb{P}[X_{\omega}(1)=x+v\Big|X_{\omega}(0)=x]=\omega(x,v)\quad v\in V_{d}$ (9)
+$\mathbb{P}[X_{\omega}(1)=x+v\Big|X_{\omega}(0)=x]=\omega(x,v)\quad v\in V_{d}.$ (9)
 The environment process is therefore given by $K_{d}^{\operatorname{Id}}\equiv K_{d}$.
@@
 We redefine $\{X_{\Lambda(\omega)}(n)\}_{n\geq 0}$ in the following manner.
-$L_{\Lambda(\omega)}g(X_{\Lambda(\omega)}(n))=\mathbb{E}[\gamma_{n}L_{\omega}g(X_{\Lambda(\omega)}(n))+(1-\gamma_{1})L_{\alpha(\omega)}g(X_{\Lambda(\omega)}(n))]\quad\forall g:\mathbb{Z}^{d}\to\mathbb{R},\penalty 10000\ \gamma_{n}\sim\text{Ber}\left(\frac{1}{2}\right)$
+$L_{\Lambda(\omega)}g(X_{\Lambda(\omega)}(n))=\mathbb{E}[\gamma_{n}L_{\omega}g(X_{\Lambda(\omega)}(n))+(1-\gamma_{n})L_{\alpha(\omega)}g(X_{\Lambda(\omega)}(n))]\quad\forall g:\mathbb{Z}^{d}\to\mathbb{R},\penalty 10000\ \gamma_{n}\sim\text{Ber}\left(\frac{1}{2}\right)$
 where $L_{\Lambda(\omega)},L_{\omega},L_{\alpha(\omega)}$ kernels of $X_{\Lambda(\omega)},X_{\omega},X_{\alpha(\omega)}$ respectively and $\gamma_{n}$ is a Bernoulli Random Variable independent of the environment and position of the walk.
@@
 Lemma 9.2.
-$K_{\gamma}(\alpha.\omega,B)=K_{1-\gamma}(\omega,\alpha.B)$
+$K_{\gamma}(\alpha.\omega,B)=K_{1-\gamma}(\omega,\alpha.B)$.
 This is the control argument using Bernoulli variables as we had mentioned, and now we shall use it explicitly to demonstrate the stationarity of the local process.
@@
 $\lambda^{\prime}\Bigl[\omega\in B_{0},\,\bar{\omega}(1)\in B_{1},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,\gamma_{n}\Bigr]$
-$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\ \bar{\omega}(k+1)\in\alpha(B_{k+1}),\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,1-\gamma_{k+1},\ldots,\gamma_{n}\Bigr]$ (11)
+$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\ \bar{\omega}(k+1)\in\alpha(B_{k+1}),\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,1-\gamma_{k+1},\ldots,\gamma_{n}\Bigr].$ (11)
 For $k=0$,
 $\lambda^{\prime}\Big[\omega\in\alpha(B_{0}),\,\bar{\omega}(1)\in B_{1},\,\ldots,\,\bar{\omega}(k)\in B_{k},\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,\gamma_{n}\Bigr]$
-$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\bar{\omega}(1)\in\alpha(B_{1})\ldots,\,\bar{\omega}(k)\in B_{k}\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,1-\gamma_{1},\ldots,\gamma_{k},\ldots,\gamma_{n}\Bigr]$ (12)
+$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\bar{\omega}(1)\in\alpha(B_{1})\ldots,\,\bar{\omega}(k)\in B_{k}\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,1-\gamma_{1},\ldots,\gamma_{k},\ldots,\gamma_{n}\Bigr].$ (12)
 To see how this implies stationarity, take $B_{i}=\pi_{0}^{-1}(A_{i})$ where $A_{i}\in\mathcal{B}(d)$. Then
@@
 Proof of Lemma 9.3: We shall prove this result using induction. First for $n=1$, we note that
-$\lambda^{\prime}[\alpha(\omega)\in B_{0},\bar{\omega}(1)\in B_{1}|\gamma]=\int_{\alpha(B_{0})}K_{\gamma}(\omega,B_{1})d\lambda^{\prime}(\omega)$
+$\lambda^{\prime}[\alpha(\omega)\in B_{0},\bar{\omega}(1)\in B_{1}|\gamma]=\int_{\alpha(B_{0})}K_{\gamma}(\omega,B_{1})d\lambda^{\prime}(\omega).$
 We perform the substitution $\omega\mapsto\alpha(\omega)$. The integral reduces to
-$\int_{B_{0}}K_{\gamma}(\alpha.\omega,B_{1})d\lambda^{\prime}\circ\alpha$
+$\int_{B_{0}}K_{\gamma}(\alpha.\omega,B_{1})d\lambda^{\prime}\circ\alpha.$
 By Lemma 9.1 and Lemma 9.2, the integral reduces to
-$\int_{B_{0}}K_{1-\gamma}(\omega,\alpha.B_{1})d\lambda^{\prime}\circ\alpha=\lambda^{\prime}(\alpha(\omega)\in B_{0},\bar{\omega}(1)\in\alpha.B_{1}|1-\gamma)$
-Suppose the result is true for $n=m$. Let $n=m+1$. We verify the statement for $k=m+1$ only for simplicity as others follow from the technique of proof. For $k=m+1$, we write
-$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times\alpha.B_{m}}K_{\gamma_{m+1}}(\omega_{m},B_{m+1})d\lambda^{\prime}(\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m})$
+$\int_{B_{0}}K_{1-\gamma}(\omega,\alpha.B_{1})d\lambda^{\prime}\circ\alpha=\lambda^{\prime}(\alpha(\omega)\in B_{0},\bar{\omega}(1)\in\alpha.B_{1}|1-\gamma).$
+Suppose the result is true for $n=m$. Let $n=m+1$. We verify the statement for $k=m+1$ only for simplicity, as others follow from the technique of proof. For $k=m+1$, we write
+$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times\alpha.B_{m}}K_{\gamma_{m+1}}(\omega_{m},B_{m+1})d\lambda^{\prime}(\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m}).$
 Perform the substitution $(\omega_{m},\dots,\omega)\mapsto(\alpha.\omega_{m},\dots,\omega)$ to get
-$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times B_{m}}K_{\gamma_{m+1}}(\alpha.\omega_{m},B_{m+1})d\lambda^{\prime}(\alpha.\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m})$
+$\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times B_{m}}K_{\gamma_{m+1}}(\alpha.\omega_{m},B_{m+1})d\lambda^{\prime}(\alpha.\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m}).$
 Using Lemma 9.2, the integral reduces to
-$\int_{B_{0}\times B_{1}\times B_{2}\times\cdots\times B_{m}}K_{1-\gamma_{m+1}}\bigl(\omega_{m},\alpha(B_{m+1})\bigr)\,d\lambda^{\prime}\bigl(\alpha(\omega_{m}),\ldots,\omega\,\Bigm|\,\gamma_{1},\ldots,\gamma_{m}\bigr)$
+$\int_{B_{0}\times B_{1}\times B_{2}\times\cdots\times B_{m}}K_{1-\gamma_{m+1}}\bigl(\omega_{m},\alpha(B_{m+1})\bigr)\,d\lambda^{\prime}\bigl(\alpha(\omega_{m}),\ldots,\omega\,\Bigm|\,\gamma_{1},\ldots,\gamma_{m}\bigr).$
 Now we write the above expression as
-$\mathbb{E}\Big[\mathbb{E}\Big[\boldsymbol{1}_{\alpha.B_{m+1}}(\bar{\omega}(m+1))\Big|\bar{\omega}(m),1-\gamma_{m+1}\Big]\boldsymbol{1}_{\bar{\omega}(m)}(\alpha.B_{m})\boldsymbol{1}_{\bar{\omega}(m-1)}(B_{m-1})\dots,\boldsymbol{1}_{\omega}(B_{0})\Big|\gamma_{1},\dots,\gamma_{m}\Big]$
+$\mathbb{E}\Big[\mathbb{E}\Big[\boldsymbol{1}_{\alpha.B_{m+1}}(\bar{\omega}(m+1))\Big|\bar{\omega}(m),1-\gamma_{m+1}\Big]\boldsymbol{1}_{\bar{\omega}(m)}(\alpha.B_{m})\boldsymbol{1}_{\bar{\omega}(m-1)}(B_{m-1})\dots,\boldsymbol{1}_{\omega}(B_{0})\Big|\gamma_{1},\dots,\gamma_{m}\Big].$
 This reduces by definition to the desired form, which completes the proof.

### v4

@@
-Ergodic Theorems and Equivalence of Green’s Kernel for Random Walks in Random Environments
-Abstract
+Ergodic Theorems for Random Walks in Random Environments
+Abstract.
 We study the ergodic properties of random walks in stationary ergodic environments without uniform ellipticity under a minimal assumption.
 There are two main components in our work.
-The first step is to adopt the arguments of Lawler to first prove a uniqueness principle.
-We use a more general definition of environments using Environment Functions.
-As a corollary, we can deduce an invariance principle for balanced environments under the assumptions.
-We also use the uniqueness principle to show that any balanced, elliptic random walk must have the same transience behaviour as the simple symmetric random walk.
-The second is to transfer the results we deduce in balanced environments to general ergodic environments(under some assumptions) using a control technique to derive a measure under which the local process is stationary and ergodic.
-As a consequence of our results, we deduce the Law of Large Numbers for the Random Walk and an Invariance Principle under our assumptions.
+The first step is to adopt the arguments of Lawler, utilizing a more general definition of environments via environment functions, to deduce an invariance principle for balanced environments under some natural assumptions.
+Secondly, we transfer these results to general ergodic environments using a control technique to derive a measure under which the local process is stationary and ergodic.
+Two important consequences of our results are the Law of Large Numbers and the Invariance Principle.
+keywords:
+Random walk, Random environment, Ergodic Theorems
 1. Introduction
@@
 This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers.
-Roughly speaking, it says there is a deterministic limiting velocity of the walk.
-This has been studied in detail in $d=1$ by Solomon[10] in 1975.
+Roughly speaking, the first one says that there is a deterministic limiting velocity of the walk.
+This has been studied in detail in $d=1$ by Solomon[solomon] in 1975.
 In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
-Informally speaking, the problem regarding the Law of Large Numbers in RWRE is the following:
+Informally speaking, the problem regarding the Law of Large Numbers in RWRE, for stationary ergodic environments, is the following:
 Conjecture 1.1.
 Does there exist a deterministic vector $v$ such that
-$\frac{X_{n}}{n}\to v\quad\text{a.e.}$
-for almost every choice of environment?
-Varadhan in [11] showed that the sequence $X_{n}/n$ has at most two deterministic limit points $v_{1},v_{2}$ and if $v_{1}\neq v_{2}$ then there is a constant $a\geq 0$ such that $v_{1}=-av_{2}$.
-Furthermore, Noam Berger in [1] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0.
-These results assume that the environment is uniformly elliptic and i.i.d..
+$\frac{X_{n}}{n}\to v$
+for almost every realisation of the environment?
+Varadhan in [v] showed that the sequence $X_{n}/n$ has at most two deterministic limit points $v_{1},v_{2}$ and if $v_{1}\neq v_{2}$ then there is a constant $a\geq 0$ such that $v_{1}=-av_{2}$.
+Furthermore, Noam Berger in [ber] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0.
+These results assume that the environment is uniformly elliptic and i.i.d.
 In our present discussion, we answer 1.1 in the affirmative for stationary ergodic nearest neighbour RWRE and show the existence of the limiting velocity under some assumptions.
+The 0-1 conjecture (which is a famous one in RWRE theory), first stated (for two dimensions) by Stephen Kalikow in [kalikow1981generalized] for uniformly elliptic i.i.d or more generally for uniformly elliptic ergodic environments, says that given a direction, the walk either escapes off or does not for almost all realisations of the environment.
+More formally
+Conjecture 1.2.
+Let $\ell\in S^{d-1}$. Let $A_{\ell}=\{X_{n}.\ell\to+\infty\}$. Then
+$\mathbb{P}[A_{\ell}]\in\{0,1\}$
+Prior work on this problem, notably in two dimensions, by Martin Zerner and Franz Merkl [merkl2001zero], has solved this for i.i.d.
+environments and has counterexamples for elliptic, ergodic environments.
+Further counterexamples exist in $d\geq 3$ in [bramson2006shortest] for stationary environments under a polynomial mixing condition, which is not equivalent to ergodic environments (it is not difficult to find counterexamples).
+The proofs crucially rely on invalidating the law of large numbers, so we expect, in view of our results, that this conjecture is also affirmative under our assumptions, but it should require a rigorous proof, which we are unable to provide.
 We shall work under a more general definition of environments using Environment Function, which we shall introduce below.
-We shall establish under some minimal assumption that this chain has a stationary measure for balanced environments, adapting arguments from [7] for any environment function.
-This shall also provide an easy proof of Theorem 1 of [4].
-This shall be derived by first establishing a uniqueness principle for concave functions on the lattice satisfying certain properties.
-We answer the following natural question for balanced random walks, which can be thought of for now as a random walk having equal probabilities of going forward or backward along each coordinate axis.
-Conjecture 1.2.
-Is every balanced, elliptic(probability of going along every axis at every point is positive) walk in $\mathbb{Z}^{d}$ recurrent in $d=1,2$ and transient in $d\geq 3$?
-We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest.
-Using the uniqueness, we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks.
-It is important to note that this section shall not deal with the notion of a Random Environment.
-This is one of the key contributions of this paper, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
+We shall establish under some minimal assumption that this chain has a stationary measure for balanced environments, adapting arguments from [Lawler1982] for any environment function.
+This shall also provide an easier and more natural proof of Theorem 1 of [guo2012quenched].
+This shall be derived by adapting the arguments from [Lawler1982].
 Our other main contribution will be to deduce the ergodicity of a certain stochastic process, which we will call the local process.
-This process chain is a projection at the origin of the “Environment Viewed from the Particle” Markov Chain or, as we will call it in this paper, the environment process with a suitable environment function.
-We consider the local process as environment process does not always possess suitable ergodic properties.
+This process is the projection at the origin of the “Environment Viewed from the Particle” Markov Chain, or, as we will call it in this paper, the environment process with a suitable environment function.
+We consider this process since it is sufficient for our purposes of deriving the law of large numbers and also in view of the fact that the environment process does not always possess suitable ergodic properties.
 This can be noted for Random Walks in Dirichlet Environments as studied by Christopher Sabot in [11].
 Our definition of environments allows us to transfer our results from the balanced case to the general environment by coupling the random walk and a reflected walk.
-We then combine a novel technique, which we shall name as the Action of Linear Transformations, and another control technique to execute a proof.
+We then combine a novel technique, which we shall name as the Action of Linear Transformations, and another control technique to execute a proof of the law of large numbers.
 As a consequence of our ergodic theorem, we shall also be able to deduce an invariance principle for general environments.
-To the best of our knowledge, these theorems are novel.
-Our results hold for all uniformly elliptic RWRE in dimensions equal to or greater than one, and the conditions we impose study the extent to which the condition of uniform ellipticity can be dropped.
+Our results hold for all uniformly elliptic RWRE in dimensions equal to or greater than one, and the conditions we impose show the extent to which the condition of uniform ellipticity can be dropped.
 Acknowledgements:
@@
 2. Prerequisites
-Before stating the main results, we formally define the setting we work under. We adopt the formalism as in [7].
+Before stating the main results, we formally define the setting we work under. We adopt the formalism as in [Lawler1982].
 Define by $\mathbb{Z}^{d}$ the integer lattice and $\{e_{i}\}_{i=1,\dots,d}$ the unit vectors. Define the set
@@
 $S(d)$ $=\{(p_{0},p_{1},\dots,p_{2d}):p_{i}\geq 0,\ p_{0}+\sum_{i=1}^{2d}p_{i}=1\}$
-$\mathcal{S}(d)$ $=\{\text{set of functions from }\mathbb{Z}^{d}\to S(d)\}.$
-Let $\mathcal{B}(d)$, $\mathcal{B}(d)^{\prime}$ be the Borel-$\sigma$ algebra on $S(d)$, $\mathcal{S}(d)$ respectively.
-Take $\mathscr{E}:\mathcal{S}(d)\to\mathcal{S}(d)$ such that $\mathscr{E}(\omega)(x)\equiv\mathscr{E}(\omega(x)):S(d)\to S(d)$ is a measurable function on $(S(d),\mathcal{B}(d))$.
-We call this the environment function.
+$\mathscr{S}(d)$ $=\{\text{set of functions from }\mathbb{Z}^{d}\to S(d)\}.$
+Let $\mathcal{B}(d)$, $\mathcal{B}(d)^{\prime}$ be the Borel-$\sigma$ algebra on $S(d)$, $\mathscr{S}(d)$ respectively.
+Consider a measurable function $\mathscr{E}:S(d)\to S(d)$. Call the environment function, the natural extension of this function on $\mathscr{S}(d)$ given by
+$\mathscr{E}(\omega)(x)=\mathscr{E}(\omega(x)).$
+We shall call such functions the environment function.
 This function essentially contains the information on the transition probabilities of the random walk based on the function $\omega$.
@@
 Definition 2.1 (Environment).
-Given an environment function $\mathscr{E}$, $\omega\in\mathcal{S}(d)$, we call the pair $(\omega,\mathscr{E})$ as the environment.
-Let $\mu$ be a probability measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. We call the environment function balanced if
+Given an environment function $\mathscr{E}$, $\omega\in\mathscr{S}(d)$, we call the pair $(\omega,\mathscr{E})$ as the environment.
+Let $\mu$ be a probability measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$. We call the environment function balanced if
 $\forall x\in\mathbb{Z}^{d},\quad\mathscr{E}_{i}(\omega)(x)=\mathscr{E}_{i+d}(\omega)(x)\quad i=1,2,\dots,d.$
@@
 $\mu[\mathscr{E}_{k}(\omega)(x)>0:\forall x\in\mathbb{Z}^{d},\forall k]=1.$
-Define the group of shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$ on $\omega\in\mathcal{S}(d)$ given by
+Define the action of the group of shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$ on $\omega\in\mathscr{S}(d)$ given by
 $(\tau_{x}\omega)(y)=\omega(x+y),\quad\forall x,y\in\mathbb{Z}^{d}.$
@@
 Notation 1.
-For $\mathscr{E}(\omega)\in\mathcal{S}(d)$, $v\in V_{d}$, $\mathscr{E}$ an environment function
+For $\mathscr{E}(\omega)\in\mathscr{S}(d)$, $v\in V_{d}$, $\mathscr{E}$ an environment function
 $\mathscr{E}(\omega)(x,v)$ $=\mathscr{E}(x,v)=\mathscr{E}_{i}(\omega)(x)\quad\text{if}\penalty 10000\ v=e_{i}$
@@
 One could think of this as independently assigning a point from $S(d)$ at every site.
-Fix $\omega\in\mathcal{S}(d)$ and an environment function $\mathscr{E}$.
+Fix $\omega\in\mathscr{S}(d)$ and an environment function $\mathscr{E}$.
 Define the Markov Chain of the Random Walk $\{X_{\mathscr{E}(\omega)}(k)\}_{k\geq 0}$ in $\mathbb{Z}^{d}$ by the transition probabilities,
@@
 respectively for every $g:\mathbb{Z}^{d}\to\mathbb{R}$.
-Define the environment process as the Markov Chain given by $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with the transition kernel given by
+Define the environment process as the Markov Chain given by $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ with the transition kernel given by
 $K_{d}^{\mathscr{E}}(\omega,B)=\sum_{v\in V_{d}}{\mathscr{E}(\omega)}(o,v)\delta_{\tau_{v}\omega}(B)\quad\forall B\in\mathcal{B}(d)^{\prime}.$
@@
 $\pi_{E}(\omega)=\{\omega(x):x\in E\}.$
-Then we define the local process associated with any discrete time Markov Chain $\{Y_{k}\}_{k\geq 0}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ as $\{\pi_{0}(Y_{k})\}_{k\geq 0}$.
 Denote by $\mathbb{P}^{\mathscr{E}({\omega})}_{x}$: the law on path of the random walk started at $x$ in the environment $(\omega,\mathscr{E})$.
 Precisely, $\mathbb{P}^{{\mathscr{E}(\omega)}}_{x}$ is a measure on $((\mathbb{Z}^{d})^{\mathbb{N}},\mathcal{G)}$ where $\mathcal{G}$ is the $\sigma$-algebra generated by cylinder sets.
-Therefore if ${\mathscr{E}(\omega)}\in\mathcal{S}(d)$ one has that
-$\mathbb{P}^{{\mathscr{E}(\omega)}}_{x}(G):\mathcal{S}(d)\to[0,1]$
+Therefore if ${\mathscr{E}(\omega)}\in\mathscr{S}(d)$ one has that
+$\mathbb{P}^{{\mathscr{E}(\omega)}}_{x}(G):\mathscr{S}(d)\to[0,1]$
 is measurable $\forall G\in\mathcal{G}$.
-Hence, we can define the annealed law to be the probability measure on $(\mathcal{S}(d)\times(\mathbb{Z}^{d})^{\mathbb{N}},\mathcal{B}(d)^{\prime}\times\mathcal{G)}$ as
+Hence, we can define the annealed law to be the probability measure on $(\mathscr{S}(d)\times(\mathbb{Z}^{d})^{\mathbb{N}},\mathcal{B}(d)^{\prime}\times\mathcal{G)}$ as
 $\mathbb{P}^{ann}_{x}(F\times G)=\int_{F}\mathbb{P}_{x}^{{\mathscr{E}(\omega)}}(G)d\mu(\omega)\quad\forall F\in\mathcal{B}(d)^{\prime},G\in\mathcal{G}.$
@@
 $d_{\mathscr{E}}(x,\omega)=E[X_{{\mathscr{E}(\omega)}}(1)-x|X_{{\mathscr{E}(\omega)}}(0)=x].$
-We make the following assumptions on the environment that is
+We make the following assumptions about the environment that is
 Assumption 2.2.
-Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. Define the quantity $c(\mathscr{E},x)$ as follows
+Let $\mu$ be a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$.
+Since we aim to drop uniform ellipticity, i.e., all transition probabilities bounded below by some constant positive value, we define the quantity $c(\mathscr{E},x)$(we may drop $\omega$ from the notation) as follows
 $c(\mathscr{E},x)=\Big[\prod_{v\in V_{d}\setminus\{0\}}\mathscr{E}(\omega)(x,v)\Big]^{\frac{1}{2d}}.$
 Then we assume that
-$\int_{\mathcal{S}(d)}c^{-p}d\mu<\infty,$
+$\int_{\mathscr{S}(d)}c^{-p}d\mu<\infty,$
 for some $p>d$.
@@
 $|x|_{1}=\sum_{j=1}^{k}|x|_{j}.$
-We use the notion of concave functions defined on the compact $L_{1}$ ball $B_{1}[0,r]\cap\mathbb{Z}^{d}$. A function $z:B_{1}[0,r]\cap\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ $(r>0)$ is said to be concave if
+We use the notion of concave functions defined on the compact $L_{1}$ ball $B_{1}[0,r]\cap\mathbb{Z}^{d}$.
+A function $z:B_{1}[0,r]\cap\mathbb{Z}^{d}\to\mathbb{R}_{\geq 0}$ $(r>0)$ is said to be concave if
 $z(x+v)+z(x-v)-2z(x)\leq 0\quad x\in\operatorname{int}B_{1}[0,r]\cap\mathbb{Z}^{d}\quad\forall v\in V_{d}\setminus\{0\},$
 where $\operatorname{int}$ denotes the topological interior of a set.
+We also define an operator on concave functions, which we shall call the Monge-Ampère Operator. For a concave function $z$ we define the operator $M$ as
+$Mz(x)=\prod_{i=1}^{d}[z(x+e_{i})+z(x-e_{i})-2z(x)].$
 Also, note that we use the notion of mutual absolute continuity of two measures.
@@
 3. Main Results and Discussions
-We first state the result that allows us to effectively deal with environments beyond uniform ellipticity. We shall demonstrate this through another result later in this section.
+We first state a result on balanced environments. This result forms the backbone for our control argument as we shall see.
 Theorem 3.1.
-Let $d\geq 1$. Assume $\mu$ is a stationary, ergodic Law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a measurable balanced environment function satisfying assumptions 2.2 and 2.3.
-Then there exists a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ such that the environment process $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ is stationary, ergodic under this measure and is mutually absolutely continuous to $\mu$.
+Let $d\geq 1$. Assume $\mu$ is a stationary, ergodic Law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a measurable balanced environment function satisfying assumptions 2.2 and 2.3.
+Then there exists a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ such that the environment process $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ is stationary, ergodic under this measure and is mutually absolutely continuous to $\mu$.
 An important corollary of this result, which we shall also state but is not directly related to our discussion, is
@@
 Let $d\geq 1$.
-Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ and $\mathscr{E}$ a measurable balanced environment function.
-Fix $\omega\in\mathcal{S}(d)$.
+Let $\mu$ be a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ and $\mathscr{E}$ a measurable balanced environment function.
+Fix $\omega\in\mathscr{S}(d)$.
 Then under assumptions 2.2 and 2.3,
@@
 converges weakly to the standard Brownian Motion on $\mathbb{R}^{d}$ with non degenerate covariance matrix of the form $(b_{i}\delta_{ij})_{1\leq i,j\leq d}$ for $\omega$ $\mu$-a.e.
-These results notably weaken the assumptions of ellipticity and derive an alternate proof of Theorem 1 of [4] for $d\geq 1$, which states that
-Theorem 3.3 (Theorem 1(i) of [4]).
+These results notably weaken the assumptions of ellipticity and derive an alternate proof of Theorem 1 of [guo2012quenched] for $d\geq 1$, which states that
+Theorem 3.3 (Theorem 1(i) of [guo2012quenched]).
 Let
-$c(x)=\prod_{i=1}^{d}[\omega(x,e_{i})]^{\frac{1}{d}}.$
+$\varepsilon_{\omega}(x)=\prod_{i=1}^{d}[\omega(x,e_{i})]^{\frac{1}{d}}.$
 If the stationary, ergodic, and elliptic environment( $\omega(x,e_{i})=\omega(x,-e_{i})$ with $\mathscr{E}=\operatorname{Id}$) satisfies the condition
-$\mathbb{E}[c(o)^{-p}]<\infty$
+$\mathbb{E}[\varepsilon(o)^{-p}]<\infty$
 for some $p>d\geq 2$, then Corollary 3.2 holds.
-To prove the above results, we derive a uniqueness theorem for concave functions on $\mathbb{Z}^{d}$, which also gives the following theorem on the transience of balanced elliptic walks.
-The result compares the walk to the simple symmetric random walk on $\mathbb{Z}^{d}$ given by the law
-$\mathbb{P}[X_{1}=x+v\Big|X_{0}=x]=\frac{1}{2d}.$
-To state our result, we recall the so-called Green’s function for a random walk on $\mathbb{Z}^{d}$.
-Definition 3.4.
-Let $(\omega,\mathscr{E})$ be a balanced environment in $\mathbb{Z}^{d}$($d\geq 1$).
-Then the Green’s function $G^{\mathscr{E}(\omega)}:\mathbb{Z}^{d}\times\mathbb{Z}^{d}\to\mathbb{R}^{+}\cup\{0\}$ is given by
-$G^{\mathscr{E}(\omega)}(x,y)=\mathbb{E}_{x}[\sum_{j=0}^{\infty}\boldsymbol{1}_{X_{\mathscr{E}(\omega)}(j)=y}],$
-where the expectation is taken conditional on $X_{\mathscr{E}(\omega)}(0)=x$. For the simple symmetric random walk, we shall denote its Green’s function as just $G$.
-Theorem 3.5.
-Let $d\geq 1$. Let $(\omega,\mathscr{E})$ be any balanced elliptic environment. Then
-$\frac{1}{2d}G(x,x)\leq G^{\mathscr{E}(\omega)}(x,x)\leq\frac{1}{c(\mathscr{E},x)^{d}}G(x,x).$
-This shows the conclusions of Pölya’s Theorem, which states
-Theorem 3.6 (Pölya).
-The simple symmetric Random Walk on $\mathbb{Z}^{d}$ is transient in $d\geq 3$ and recurrent in $d=1,2$.
-carries on to any random walk in a balanced, elliptic environment, therefore answering Conjecture 1.2 in the affirmative.
-Recall that a random walk is transient or recurrent depending on $G^{\mathscr{E}({\omega})}(x,x)<\infty$ or $G^{\mathscr{E}({\omega})}(x,x)=\infty$.
+To prove the above results, we shall need to bound a certain resolvent using martingale arguments and bounds on the Monge-Ampère Operator using arguments from 3.1.
 To state the next result, which connects balanced and general unbalanced regimes, we need the notion of embedded environments.
@@
 3.1. Embedded Environments
-Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$.
+Let $\mu$ be a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$.
 The embedded environment function, denoted by $\Lambda\equiv\Lambda_{\omega}$, is a balanced environment function given by
-$\Lambda_{\omega}(x,v)$ $=\frac{\omega(x,v)+\omega(x,-v)}{2}$ (1)
+(1) $\Lambda_{\omega}(x,v)$ $=\frac{\omega(x,v)+\omega(x,-v)}{2}$
 for every $v\in V_{d}$.
+The embedding function basically averages the probabilities of the opposing directions when the environment function is the Identity function.
 We shall assume without loss of generality that the environment function is Identity.
@@
 Assume $d\geq 1$.
-Let $\mu$ be an stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$.
+Let $\mu$ be a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$.
 Then suppose the environment process for the Embedded Environment satisfies the conclusions of 3.1.
-Then there exists a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic.
+Then there exists a stationary ergodic measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic.
 This theorem, combined with Theorem 3.1, gives us the following that
 Theorem 3.5.
-Let $d\geq 1$, and let $\mu$ a stationary ergodic environment law on $\mathcal{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
+Let $d\geq 1$, and let $\mu$ be a stationary ergodic environment law on $\mathscr{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
   1. (i)
@@
 The Action of Linear Transformations proves crucial in performing this step.
-This might and should appear strange without prior context, but we shall see this arising naturally as an artifact of our proof technique.
 4. An outline of the Proofs
-Let us begin this section by seeing why an invariant measure for the local process guarantees answers to Conjectures 1.1. We assume for now that Theorems 3.1 and 3.7 to be true and prove 3.8
+Assume without loss of generality that $X_{\mathscr{E}(\omega)}(0)=0$ where $\mathscr{E}$ is an environment function, for the proofs unless otherwise indicated.
+Let us begin this section by seeing why an invariant measure for the local process guarantees answers to Conjectures 1.1. We assume for now that Theorems 3.1 and 3.4 are true and prove 3.5.
 4.1. Proof of Theorem 3.5:
-First note that by Theorems 3.1 and 3.7 we have a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic since Assumptions 2.2 and 2.3 are satisfied.
+First note that by Theorems 3.1 and 3.4 we have a stationary ergodic measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic since Assumptions 2.2 and 2.3 are satisfied.
 Call this measure $\mathbb{Q}$.
-Assume without loss of generality that $X_{\omega}(0)=0$.
 Proof of Conjecture 1.1: We work with the environment following the distribution $\mathbb{Q}$.
@@
 $\sum_{k=0}^{n-1}\frac{X_{\omega}(k+1)-X_{\omega}(k)-d(0,\tau_{X_{\omega}(k)}\omega)}{n}\to 0\hskip 5.0pt\mathbb{P}_{0}^{\omega}\ \ \text{a.e.}$
-And by Birkhoff’s Ergodic Theorem [2],
-$\sum_{k=0}^{n-1}\frac{d(0,\tau_{X_{\omega}(k)}\omega)}{n}\to\int_{\mathcal{S}(d)}d(0,{\omega})d\mathbb{Q}\hskip 5.0pt\ \ \mathbb{Q}\ \ \text{a.e.}$
-Combining both gives us the theorem $\mathbb{Q}_{0}=\mathbb{Q}\rtimes\mathbb{P}_{0}^{\omega}$ a.e., which by mutual absolute continuity of measures shows it holds $\mu^{0}=\mu\rtimes\mathbb{P}_{0}^{\omega}$ a.e.
+And by Birkhoff’s Ergodic Theorem [birkhoff1931proof],
+$\sum_{k=0}^{n-1}\frac{d(0,\tau_{X_{\omega}(k)}\omega)}{n}\to\int_{\mathscr{S}(d)}d(0,{\omega})d\mathbb{Q}\hskip 5.0pt\ \ \mathbb{Q}\ \ \text{a.e.}$
+Combining both gives us the theorem $\mathbb{Q}^{0}=\mathbb{Q}\rtimes\mathbb{P}_{0}^{\omega}$ a.e., which by mutual absolute continuity of measures shows that it holds $\mu^{0}=\mu\rtimes\mathbb{P}_{0}^{\omega}$ a.e.
 Proof of Theorem 3.5(ii): First we note that under the sequence of filtrations $\{\mathcal{F}_{k}\}_{k\geq 1}$,
@@
 has the entries
-$(\Sigma_{k})_{ij}=\begin{cases}T(k)_{i}+T(k)_{i+d}-(T_{i}(k)-T_{i+d}(k))^{2}&\text{if }i=j,\\ -[T_{i}(k)-T_{i+d}(k)][T_{j}(k)-T_{j+d}(k)],&\text{if }i\neq j\end{cases}$
+$(\Sigma_{k})_{ij}=\begin{cases}T_{i}(k)+T_{i+d}(k)-(T_{i}(k)-T_{i+d}(k))^{2}&\text{if }i=j,\\ -[T_{i}(k)-T_{i+d}(k)][T_{j}(k)-T_{j+d}(k)],&\text{if }i\neq j\end{cases}$
 for $1\leq i,j\leq d$.
-Therefore by 3.7,
+Therefore by 3.4,
 $\frac{1}{n}\sum_{k=0}^{n-1}\Sigma_{k}\to\Sigma\quad\omega\penalty 10000\ \mu\textit{-a.e.}$
 where $\Sigma$ is a positive definite matrix under assumption 2.3. Therefore, we conclude the theorem via the invariance principle for Martingales.
-The proof of Theorem 3.5 is postponed and will be shown as a natural consequence of our uniqueness principle.
-Thus, we shall devote the rest of the paper to proving Theorems 3.1 and Theorem 3.7.
-We divide the proof into various sections in order to properly motivate the proof.
-4.2. Discussion on Theorems 3.1, Corollary 3.2 and Theorem 3.5:
-This section provides an alternate proof of Theorem 3.3 but generalises it to the general case of stationary, ergodic environments and arbitrary environment functions.
+Thus, we shall devote the rest of the paper to proving Theorems 3.1 and Theorem 3.4. We divide the proof into various sections in order to properly motivate the proof.
+4.2. Discussion on Theorems 3.1, Corollary 3.2:
+This section provides an easier proof of Theorem 3.3 but generalises it to the general case of stationary, ergodic environments and arbitrary environment functions.
 This theorem is important for verifying the conditions of Theorem 3.4 under the assumptions of Theorem 3.5.
-The two main themes in this proof are the use of Periodised Environments as used in [7] and our uniqueness principle.
+The two main themes in this proof are the use of Periodised Environments as used in [Lawler1982] and a uniqueness principle for concave functions.
 We shall prepare the reader with the prerequisites for the proof in the next section and fully motivate the proof.
-The proof in [7] adopts the arguments of Papanicolaou and Varadhan [8] for diffusion processes with random coefficients for the random walk for uniformly elliptic balanced environments.
-Our main step is to refine the arguments in [7] for non-uniformly elliptic balanced environments under assumptions 2.2 and 2.3 to get our uniqueness principle.
-The ideas of Krylov [6] were used to estimate the solutions of the discrete Monge-Ampère Equation by Lawler.
-We adopt these ideas to bound a certain resolvent as in [7] under assumptions 2.2 and 2.3.
-We further use this uniqueness principle to derive Theorem 3.5.
-We emphasise that Theorem 3.5 is not a result on Random Walk in Random Environments but on general balanced elliptic walks.
+The proof in [Lawler1982] adopts the arguments of Papanicolaou and Varadhan [PapanicolaouVaradhan1982] for diffusion processes with random coefficients for the random walk for uniformly elliptic balanced environments.
+Our main step is to refine the arguments in [Lawler1982] for non-uniformly elliptic balanced environments under assumptions 2.2 and 2.3 to get our uniqueness principle.
+The ideas of Krylov [Krylov1971] were used to estimate the solutions of the discrete Monge-Ampère Equation by Lawler.
+We adopt these ideas to bound a certain resolvent as in [Lawler1982] under assumptions 2.2 and 2.3.
 4.3. Discussion on Theorems 3.4:
@@
 The first being the use of Action of Linear Transformations on the Environment.
-This approach is novel to the best of my knowledge.
+This approach is novel in the sense that it generalises the affine shift operator on the environment to linear transforms.
 Let us describe this theory in brief.
@@
 These notions prove extremely fundamental to our arguments as we shall see.
-We shall prepare the reader accordingly in the next section for the proof.
+We shall prepare the reader accordingly in the upcoming sections for the proof.
 The second core theme in this proof is to pass from the embedded environment to the original environment by exhibiting stationarity of the local process through a control argument using Bernoulli Random Variables Independent of the environment.
@@
 The control argument will therefore allow us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process.
-The construction is specific to the proof and shall be demonstrated accordingly during the proof.
-The main hurdle in the proof is to show stationarity of the local process. The ergodicity of the local process follows by adapting arguments from [3] suitably, which shall be demonstrated.
-5. Preparation for the Proof of Theorem 3.1, Corollary 3.2 and Theorem 3.5 :
-Fix an environment ${\omega}\in\mathcal{S}(d)$ and $\mathscr{E}$ a balanced measurable environment function on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$.
+In particular, the local process of the walk shall be equal in law to the local process of the random walk in the embedded environment.
+The control argument shall help us achieve this by sort of averaging the original and the reflected, as might be noted in the choice of the embedding function.
+The main hurdle in the proof is to show stationarity of the local process.
+The ergodicity of the local process follows by adapting arguments from [BolthausenSznitman2002] suitably, which shall be demonstrated.
+5. Preparation for the Proof of Theorem 3.1, Corollary 3.2 :
+Fix an environment ${\omega}\in\mathscr{S}(d)$ and $\mathscr{E}$ a balanced measurable environment function on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$.
 We first state our uniqueness result. We shall use this result crucially in the upcoming sections.
@@
 Define the following:
-$D_{n}=\{x\in\mathbb{Z}^{d}:|x|_{1}\leq n\}$
-$\partial D_{n}=\{x\in D:|x|_{1}=n\}$
-and
-$D_{n}^{\circ}=D\setminus\partial D_{n}.$
+$D_{n}$ $=\{x\in\mathbb{Z}^{d}:|x|_{1}\leq n\},$
+$\partial D_{n}$ $=\{x\in D_{n}:|x|_{1}=n\},$
+$D_{n}^{\circ}$ $=D_{n}\setminus\partial D_{n}.$
 Let $x\in D_{n}^{\circ}$. Define the second-order difference operators for a concave function $z:D_{n}\to\mathbb{R}$ given by
@@
 Define the class of concave functions $\mathscr{A}(\mathscr{E},f)$ containing concave functions $u(x)$ satisfying the properties
-  1. 1.
+  1. (1)
     $u\equiv 0$ on $\partial D_{n}$,
-  2. 2.
-    $L_{\mathscr{E}(\omega)}u(x)\leq-f(x)+u(x)\quad\forall x\in D_{n}^{\circ}$.
+  2. (2)
+    $|Mz(x)|^{\frac{1}{d}}\geq\frac{f(x)}{c(\mathscr{E},x)}\quad x\in D_{n}^{\circ}$
 Theorem 5.1 (Uniqueness Principle).
-Let $(\omega,\mathscr{E})$ be any balanced, elliptic environment.
-There exists an unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that $L_{\mathscr{E}(\omega)}z(x)=-f(x)+z(x)\quad\forall x\in D_{n}^{\circ}$ given by
-$z(x)=\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)})(x)],$
-where the expectation is taken assuming $X_{\mathscr{E}(\omega)}(0)=x\in D_{n}^{\circ}$. Furthermore, this function is also given alternately by
+Let $(\omega,\mathscr{E})$ be any balanced, elliptic environment. There exists a unique function $z\in\mathscr{A}(\mathscr{E},f)$ such that
+$|Mz(x)|^{\frac{1}{d}}=\frac{f(x)}{c(\mathscr{E},x)}.$
+Furthermore, this function is also given alternately by
 $z(x)=\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x),\quad\forall x\in D_{n}.$
@@
 $K^{\mathscr{E}}_{d}(g(Y_{j}))=\sum_{i=1}^{d}\mathscr{E}_{i}(Y_{j})(0)[g(\tau_{e_{i}}Y_{j})+g(\tau_{-e_{i}}Y_{j})]+\mathscr{E}_{0}(Y_{j})(0)g(Y_{j})$
-for all $g:\mathcal{S}(d)\to\mathbb{R}^{d}$. Therefore 3.2 is equivalent to the following theorem,
+for all $g:\mathscr{S}(d)\to\mathbb{R}^{d}$. Therefore 3.2 is equivalent to the following theorem, which is a restatement of Theorem 3.1,
 Theorem 5.3.
-Let $\mu$ be a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$.
+Let $\mu$ be a stationary ergodic measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$.
 Then under assumptions 2.2 and 2.3, there exists a stationary ergodic measure $\lambda$ mutually absolutely continuous to $\mu$ under which $\{Y_{j}\}$ is stationary and ergodic.
-To prove stationarity of the chain is enough to conclude ergodicity via arguments in [3]. Under 5.3,
-$b=\int_{\mathcal{S}(d)}\mathscr{E(}\omega)d\lambda.$
+To prove stationarity of the chain is enough to conclude ergodicity via arguments in [BolthausenSznitman2002]. Under 5.3,
+$b=\int_{\mathscr{S}(d)}\mathscr{E(}\omega)d\lambda.$
 We shall work towards proving Theorem 5.3 under assumptions 2.2 and 2.3. We shall now introduce the notion of Periodised Environments as mentioned in section 4.2, fundamental to the proof of 5.3.
@@
 $(x_{1},x_{2},\dots,x_{d})\sim(y_{1},y_{2},\dots,y_{d})\iff\frac{(x_{i}-y_{i})}{2n}\in\mathbb{Z}\penalty 10000\ \text{for each}\penalty 10000\ i=1,2,\dots,d.$
-Let $\mathbb{Z}^{d}/\sim=T_{n}$. Clearly $|T_{n}|=(2n)^{d}$. We shall also identify $T_{n}$ as the set in $\mathbb{Z}^{d}$ given by
+Let $\mathbb{Z}^{d}/\sim\penalty 10000\ =T_{n}$. Clearly $|T_{n}|=(2n)^{d}$. We shall also identify $T_{n}$ as the set in $\mathbb{Z}^{d}$ given by
 $T_{n}=\{-n+1,-n+2,\dots,0,1,2,\dots,n\}^{d}.$
 Definition 5.4 (Periodic Environment).
-We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$. Let $\mathcal{S}_{n}$ denote the set of all such $\omega$.
-Fix $\omega\in\mathcal{S}(d)$. We take the periodic environment $(\omega_{n},\mathscr{E})$ as
+We call $\omega$ a periodic environment if $\omega$ is a function $\omega:T_{n}\to S(d)$. Let $\mathscr{S}_{n}$ denote the set of all such $\omega$.
+Fix $\omega\in\mathscr{S}(d)$. We take the periodic environment $(\omega_{n},\mathscr{E})$ as
 $\omega_{n}(x)$ $=\omega([x]_{\sim})$
@@
 $g(x)=g([x]_{\sim}).$
-Denote by $R_{n}^{\omega}$ the resolvent given by
-$R_{n}^{\omega}g(x)=\sum_{j=0}^{\infty}(1-\frac{1}{n^{2}})^{j}L_{\mathscr{E}({\omega})}^{j}g(x)$
+For an environment $\omega$, denote by $R_{n}^{\omega}$ the resolvent given by
+$R_{n}^{\omega}g(x)=\sum_{j=0}^{\infty}(1-\frac{1}{n^{2}})^{j}L_{\mathscr{E}({\omega_{n}})}^{j}g(x)$
 For $g:T_{n}\to\mathbb{R}$, define the $l_{p}$($1\leq p\leq\infty)$ norms with respect to the normalized counting measure on $T_{n}$ by
@@
 […]
-As in the original paper by [7], in the following lemma we state a bound on the resolvent which in turn proves 5.3.
+As in the original paper by [Lawler1982], in the following lemma we state a bound on the resolvent which in turn proves 5.3.
 Lemma 5.5.
-There exists $c_{1}\equiv c_{1}(d)>0$ such that $\forall\omega\in\mathcal{S}_{n}$,
+There exists $c_{1}\equiv c_{1}(d)>0$ such that $\forall\omega\in\mathscr{S}_{n}$,
 […]
-To show that 5.5 $\implies$ 5.3, we first state the so called multidimensional ergodic theorem[Chapter 6, Theorem 2.8 of [5]]
+To show that 5.5 $\implies$ 5.3, we first state the so called multidimensional ergodic theorem[Chapter 6, Theorem 2.8 of [Krengel1985]].
 Lemma 5.6.
@@
 Then we have:
-  1. 1.
+  1. (1)
     $\mu_{n}\to\mu$ weakly $\mu$-a.e.
-  2. 2.
+  2. (2)
     For every measurable $f:S(d)\to\mathbb{R}$ such that $f\in L^{p}(\mu)$ where $p\geq 1$, $\mu_{n}f\to\mu f$ $\mu$-a.e.
@@
 Note that $\phi_{n}$ is the density with respect to the normalised counting measure on $T_{n}$.
-Therefore, one can check (or refer to Section 2 of [4]) the invariant measure for $\{\omega_{n}(\hat{X}_{k})\}_{k\geq 0}$ is given by
-$\lambda_{n}(\tau_{x}\omega_{n})=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
-Since $\mathcal{S}(d)$ is compact with respect to the product topology, there exists a subsequence $n_{k}$ such that
+Therefore, one can check (or refer to Section 2 of [guo2012quenched]) the invariant measure for $\{\omega_{n}(\hat{X}_{k})\}_{k\geq 0}$ is given by
+$\lambda_{n}=\sum_{x\in T_{n}}\delta_{\tau_{x}\omega_{n}}\frac{\phi_{n}(x)}{(2n)^{d}}.$
+Since $\mathscr{S}(d)$ is compact with respect to the product topology, there exists a subsequence $n_{k}$ such that
 $\lambda_{n_{k}}\to\lambda\penalty 10000\ \text{weakly}$
-for some measure $\lambda$ on $\mathcal{S}(d)$ and $K^{\mathscr{E}}_{d}$ is invariant under $\lambda$.
-Now observe that
+for some measure $\lambda$ on $\mathscr{S}(d)$ and $K^{\mathscr{E}}_{d}$ is invariant under $\lambda$. Now observe that
 […]
@@
 This shows
-$\frac{d\lambda_{n}}{d\mu_{n}}||_{\frac{p}{p-1}}<\infty\quad\mu\text{-a.e.}$
-Since $\lambda_{n}$ extends as a measurable function over the entire $\mathcal{S}(d)$, we have that the family of functions
+[…]
+for some constant $C$.
+Since $\lambda_{n}$ extends as a measurable function over the entire $\mathscr{S}(d)$, we have that the family of functions
 $\big\{\frac{d\lambda_{n}}{d\mu_{n}}\big\}_{n\geq 1}$
 is uniformly integrable with respect to $\mu$. Therefore by Dunford-Pettis Theorem, $\lambda\ll\mu$ and
-$\int_{\mathcal{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<\infty.$
-Referring to arguments from [3], $\lambda$ is ergodic and $\lambda\ll\mu$ and $\lambda\gg\mu$. This concludes the proof.
+$\int_{\mathscr{S}(d)}|\frac{d\lambda}{d\mu}|^{\frac{p}{p-1}}d\mu<C^{\frac{p}{p-1}}.$
+Referring to arguments from [BolthausenSznitman2002], $\lambda$ is ergodic and $\lambda\ll\mu$ and $\lambda\gg\mu$. This concludes the proof.
 Thus, in view of the arguments in this section, we shall prove Lemma 5.5 in the next section.
@@
 We assume the Uniqueness Principle[Theorem 5.1] to be true. We first deduce Lemma 5.5 under this assumption and then prove Theorem 5.1 in the next section.
-Proof of Lemma 5.5 : Let $\omega\in\mathcal{S}_{n}$ and let $f:D_{n}\to[0,\infty)$. Let
+Proof of Lemma 5.5 : Let $\omega\in\mathscr{S}_{n}$ and let $f:D_{n}\to[0,\infty)$. Let
 $\tau_{n}=\inf\{m:X_{\mathscr{E}(\omega)}(m)\in\partial D_{n}\}$
@@
 […]
-To see the equivalence, one notes the fact that $z(x)=Qf(x)$.
-Proof of Lemma 6.2: To prove the Lemma, we consider the Monge-Ampère Operator for a given concave function $z$ given by
+Proof of Lemma 6.2: First we note the equivalence stated above. Recall the Monge-Ampère Operator for a given concave function $z$ given by
 $Mz(x)=\prod_{i=1}^{d}\Delta_{i}z(x).$
 Note that from Theorem 5.1 one has that
-$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)=f(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x).$
-Define the sets for $x\in D_{n}^{\circ}$
+$\sum_{i=1}^{d}-\mathscr{E}(x,e_{i})\Delta_{i}z(x)\geq|Mz|^{\frac{1}{d}}c(\mathscr{E},x)=f(x)$
+Therefore, one can show that for every $j\geq 1$,
+$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]\leq\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\mathscr{E}(\omega)}(k)).$
+[…]
+Define the sets for $x\in D_{n}^{\circ}$,
 $I(x)=\{a\in\mathbb{R}^{d}:z(x+e_{i})-z(x)\leq a_{i}\leq z(x)-z(x-e_{i}),\text{for}\penalty 10000\ i=1,2,\dots,d\}.$
 Observe that
-$meas(I(x))\leq\frac{f(x)^{d}}{c(\mathscr{E},x)^{d}},$
-where $meas(\cdot)$ is the Lebesgue measure on $\mathbb{R}^{d}$.
-Now we state an easily provable lemma as in [7].
-Lemma 6.3 (Lemma 8 of [7]).
+$meas(I(x))=|Mz(x)|=\frac{f(x)^{d}}{c(\mathscr{E},x)^{d}},$
+where $meas(\cdot)$ is the Lebesgue measure on $\mathbb{R}^{d}$. Now we state an easily provable lemma as in [Lawler1982].
+Lemma 6.3 (Lemma 8 of [Lawler1982]).
 Let $a\in\mathbb{R}^{d}.$ and let $r(x)=a.x+b$ for $b>0$. If $r(x)\geq z(x)$ for every $x\in D_{n}$ and $\exists x_{0}\in D_{n}^{\circ}$ such that $r(x_{0})=z(x_{0})$, then $a\in I(x_{0})$.
@@
 […]
-Therefore
 $A\subset\bigcup_{x\in D_{n}^{\circ}}I(x),$
@@
 We now bound the resolvent for non-negative $g:T_{n}\to\mathbb{R}$(and extended as mentioned before) as follows,
-$\mathbb{E}_{x}\big[\sum_{j=0}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega)}(j)\big)]\leq\mathbb{E}_{x}\big[\sum_{j=0}^{\tau_{Kn}}g(X_{\mathscr{E}(\omega)}(j))\big]+\mathbb{E}_{x}\big[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega)}(j)\big)],$
-where $x\in T_{n}$.
-Using 6.1, one has that
+$\mathbb{E}_{x}\big[\sum_{j=0}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega_{n})}(j)\big)]\leq\mathbb{E}_{x}\big[\sum_{j=0}^{\tau_{Kn}}g(X_{\mathscr{E}(\omega_{n})}(j))\big]+\mathbb{E}_{x}\big[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g(X_{\mathscr{E}(\omega_{n})}(j)\big)],$
+where $x\in T_{n}$. Using 6.1, one has that
 […]
-where $c^{\prime}\equiv c^{\prime}(d)$.
-The last inequality follows due to the periodic nature of $g$.
+where $c^{\prime}\equiv c^{\prime}(d)$. The last inequality follows due to the periodic nature of $g$.
 For the other summand, we write it as follows,
-$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\Bigg]$ $=\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}>n^{2}\}}\Bigg]$ (2)
-$\quad+\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg].$
+$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\Bigg]$ $=\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\mathbf{1}_{\{\tau_{Kn}>n^{2}\}}\Bigg]$
+$\quad+\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg].$
 The first sum can be bounded as follows,
@@
 For the other summand, observe that
-$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}.$
+$\mathbb{E}_{x}\Bigg[\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\mathbf{1}_{\{\tau_{Kn}\leq n^{2}\}}\Bigg]\leq\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]\cdot\|R_{n}g\|_{\infty}.$
 The last line follows from the Strong Markov property since
 $\{\tau_{Kn}\leq n^{2}\}\in\mathcal{F}_{\tau_{Kn}}=\{A:\forall\,t\in\mathbb{N},\,\{\tau_{Kn}\leq t\}\cap A\in\mathcal{F}_{t+1}\}$
-$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega)}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right).$
-Since $|X_{\mathscr{E}(\omega)}(n)|_{1}^{2}$ is a submartingale, we have by Doob’s inequality that
-$\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]=\mathbb{P}_{x}[\max_{0\leq m\leq n^{2}}|X_{{\mathscr{E}(\omega)}}(m)|_{1}^{2}\geq K^{2}n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}.$
+$\sum_{j=\tau_{Kn}+1}^{\infty}\left(1-\frac{1}{n^{2}}\right)^{j}g\!\left(X_{\mathscr{E}(\omega_{n})}(j)\right)\in\mathcal{F}^{\prime}_{\tau_{Kn}}=\sigma\!\left(\{X_{\mathscr{E}(\omega)}(k):k>\tau_{Kn}\}\right).$
+Since $\{|X_{\mathscr{E}(\omega)}(k)|_{1}^{2}\}_{k\geq 1}$ is a submartingale, we have by Doob’s inequality that
+$\mathbb{P}_{x}[\tau_{Kn}\leq n^{2}]=\mathbb{P}_{x}[\max_{0\leq m\leq n^{2}}|X_{{\mathscr{E}(\omega_{n})}}(m)|_{1}^{2}\geq K^{2}n^{2}]\leq\frac{n^{2}+|x|_{1}^{2}}{K^{2}n^{2}}.$
 Furthermore, observe that since $x\in T_{n}$, we must have $|x|^{2}_{1}\leq n^{2}d^{2}$ for every $x\in T_{n}$. Therefore, we have
@@
 By construction, the assumptions of Lemma 5.5 are satisfied, thereby completing the proof.
-Therefore, it is now enough to prove Theorem 5.1. We shall in the next section also see how this leads to a natural proof of Theorem 3.5.
-7. Proof of Theorem 5.1 and Theorem 3.5
+Therefore, it is now enough to prove Theorem 5.1.
+7. Proof of Theorem 5.1
 Proof of Theorem 5.1: We begin by noting that the class $\mathscr{A}(\mathscr{E},f)$ is non-empty. Indeed, consider
@@
 We claim that this function satisfies the second condition of the Theorem. Indeed if it does not then $\exists x_{0}\in D_{n}^{\circ}$ such that
-$L_{\mathscr{E}(\omega)}z(x_{0})<-f(x_{0})+z(x_{0}).$
+$|Mz|^{\frac{1}{d}}>\frac{f(x_{0})}{c(\mathscr{E},x_{0})}\geq 0$
 One can choose $0<\beta<z(x_{0})$ such that
-$\sum_{i=1}^{d}\mathscr{E}(\omega)(x_{0},e_{i})[z(x_{0}+e_{i})+z(x_{0}-e_{i})-2\beta]<-f(x_{0}).$
-It is again easy to show that
+$\Big[\prod_{i=1}^{d}(2\beta-z(x+e_{i})-z(x-e_{i}))\Big]^{\frac{1}{d}}>\frac{f(x_{0})}{c(\mathscr{E},x_{0})}$
+It is again easy to show that,
 $z^{\prime}(x)=\begin{cases}\beta,&\text{if }x=x_{0},\\ z(x),&\text{otherwise}.\end{cases}$
 is a member of $\mathscr{A}(\mathscr{E},f)$. This contradicts the choice of $z(x)$.
-One also notes that for every $j\geq 1$,
-$\mathbb{E}_{x}[z(X_{\mathscr{E}(\omega)}(j\wedge\tau_{n}))-z(X_{\mathscr{E}(\omega)}(0))]=\sum_{k=0}^{(j-1)\wedge\tau_{n}}-f(X_{\mathscr{E}(\omega)}(k)).$
-Letting $k\to\infty$ completes the proof.
-Proof of Theorem 3.5: Fix any balanced elliptic environment $(\omega,\mathscr{E})$.
-Let $f:D_{n}\to[0,\infty)$ be a function.
-Observe that for every concave function $u\in\mathscr{A}(\mathscr{E},f)$ one has
-$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\leq\frac{c(\mathscr{E},x)^{d}}{2d}\sum_{i=1}^{d}\Delta_{i}u(x).$
-Denote by $i$ the environment function which gives $\frac{1}{2d}$ on all the coordinates except the first one, where it assigns 0.
-Therefore, we have that
-$\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)\subset\mathscr{A}(\mathscr{E},f).$
-Denote by $\{X_{j}\}_{j\geq 0}$ the simple symmetric random walk. Then we have by Theorem 5.1 that
-$\inf_{u\in\mathscr{A}\left(i,\frac{f(x)}{c(\mathscr{E},x)^{d}}\right)}u(x)\geq\inf_{u\in\mathscr{A}(\mathscr{E},f)}u(x).$
-or by using the other form,
-$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}]\geq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))].$
-Similarly, we note that
-$\sum_{i=1}^{d}\mathscr{E}(x,e_{i})\Delta_{i}u(x)\geq\sum_{i=1}^{d}\Delta_{i}u(x).$
-This shows that
-$\mathscr{A}(\mathscr{E},f)\subset\mathscr{A}\left(i,\frac{f(x)}{2d}\right).$
-As above, one can show that
-$\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\geq\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})].$
-Therefore, we have established the following result.
-Theorem 7.1.
-Let $f:D_{n}\to[0,\infty)$ be a function. Let $(\omega,\mathscr{E})$ be a balanced elliptic environment. Then for all $x\in D_{n}^{\circ}$, we have the inequality,
-$\frac{1}{2d}\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{j})]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}f(X_{\mathscr{E}(\omega)}(j))]\leq\mathbb{E}_{x}[\sum_{j=0}^{\tau_{n}}\frac{f(X_{j})}{c(\mathscr{E},X_{j})^{d}}].$
-One can now conclude Theorem 3.5 by taking the function $f:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$
-$f(u)=\boldsymbol{1}_{x}(u)$
-and letting $n\to\infty$.
-8. Preparation for the proof of Theorem 3.7:
+8. Preparation for the proofs of Theorem 3.4:
 We begin this section by discussing the theory of Action of Linear Transformations.
@@
 8.1. Action of Linear Transformations:
-Fix $\omega\in\mathcal{S}(d)$.
-For this proof For a given linear transformation $T\in\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}^{d},\mathbb{Z}^{d})$, the action of $T$ on $\omega=\{\omega(x,v)\}_{x\in\mathbb{Z}^{d},v\in V_{d}}$ is given by
+Fix $\omega\in\mathscr{S}(d)$.
+For a given linear transformation $T\in\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}^{d},\mathbb{Z}^{d})$, the action of $T$ on $\omega=\{\omega(x,v)\}_{x\in\mathbb{Z}^{d},v\in V_{d}}$ is given by
 $T.\omega(x,v):=T(\omega)(x,v)=\omega(Tx,v).$
@@
 The induced random walk under the transformation $T$ in $\mathbb{Z}^{d}$ started from $x\in\mathbb{Z}^{d}$ is given by
-$\mathbb{P}_{0}^{\omega}(T(X_{n})=T(X_{n-1})+T(e_{i}))=\mathbb{P}_{0}^{\omega}(X_{n}=X_{n-1}+e_{i})=\omega(X_{n-1},e_{i}).$
+$\mathbb{P}_{0}^{\omega}(T(X_{\omega}(n))=T(X_{\omega}({n-1}))+T(e_{i}))=\mathbb{P}_{0}^{\omega}(X_{\omega}(n)=X_{\omega}({n-1})+e_{i})=\omega(X_{n-1},e_{i}).$
 Call the transition kernel of the environment under this new random walk, i.e., the process $\tau_{T(X_{\omega}(n))}\omega$, $K^{T}_{d}(\omega,B)$. Then we have the result,
@@
 Corollary 8.3.
-If $\tau_{X_{\omega}(n)}\omega$ has a stationary ergodic measure $Q$, $\tau_{T(X_{\omega}(n))}\omega$ is also stationary and ergodic under $Q\circ T$.
+If $\tau_{X_{\omega}(n)}\omega$ has a stationary measure $Q$, $\tau_{T(X_{\omega}(n))}\omega$ is also stationary under $Q\circ T$.
 Proof.
-$\int_{\mathcal{S}(d)}K_{d}^{T}(\omega,B)dQ\circ T=\int_{\mathcal{S}(d)}K_{d}(T.\omega,T(B))dQ\circ T=Q\circ T(B)\quad\forall B\in\mathcal{B}({d})^{\prime}$
+$\int_{\mathscr{S}(d)}K_{d}^{T}(\omega,B)dQ\circ T=\int_{\mathscr{S}(d)}K_{d}(T.\omega,T(B))dQ\circ T=Q\circ T(B)\quad\forall B\in\mathcal{B}({d})^{\prime}$
 We shall use these notions fundamentally in our proof.
 We also note that showing stationarity of the local process is sufficient to exhibit its ergodicity, which we state in the following lemma.
-We shall prove the lemma by adapting arguments from [3].
+We shall prove the lemma by arguments from [BolthausenSznitman2002].
 Lemma 8.4.
-If $\mathbb{Q}$ is a measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary then it is also ergodic.
-Let $\mathcal{S}^{\prime}=S(d)^{\mathbb{N}}$ denote all $S(d)$ valued trajectories of the local process.
-Let $T$ denote the canonical shifts or the $n^{th}$ iterate of the process.
+Let $\mathbb{Q}$ be a measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}$ is stationary.
+Further assume $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}=\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ in distribution under $\mathbb{Q}$.
+Then $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic under $\mathbb{Q}$.
+Proof.
+Stationarity follows trivially from the assumptions. We just have to show the ergodicity of the local process.
+Let $\mathscr{S}^{\prime}=S(d)^{\mathbb{N}}$ denote all $S(d)$ valued trajectories of the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$.
+Let $\theta$ denote the canonical shifts or the $n^{th}$ iterate of the process.
 Denote by $\tilde{\mathcal{B}}$ the product $\sigma$-algebra.
-Let $A$ be a set which is invariant under $T$, i.e., $T^{-1}A=A$.
-Let $h(\omega)=\tilde{P}_{\omega}(A)$ where $\tilde{P_{\omega}}$ is the law if the chain starts from $\{\omega(x)\}_{x\in\mathbb{Z}^{d}}\sim\mathbb{Q}$. We must show
-$\tilde{P}_{\mathbb{Q}}(A)=\int_{\mathcal{S}(d)}\tilde{P}_{\omega}(A)d\mathbb{Q}(\omega)=0\ \ or\ \ 1.$
-Denote $\bar{\omega}(n)=\tau_{X_{\omega}(n)}\omega$.
-Start by noting that under the canonical filtrations sequence $\mathscr{G}_{k}=\sigma(\omega,\bar{\omega}(1),\dots,\bar{\omega}(k-1))$,
-$h(\bar{\omega}(n))$ is a $\tilde{P}_{\mathbb{Q}}$ martingale since
-$\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}\circ T^{n}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=h(\bar{\omega}(n))\quad\mathbb{Q}\text{-a.e.}$
+Let $A$ be a set which is invariant under $\theta$, i.e., $\theta^{-1}A=A$.
+Let $h(\omega)=\tilde{P}_{\omega}(A)$ where $\tilde{P}_{\omega}$ is the law if the chain starts from
+$\{\omega(x)\}_{x\in\mathbb{Z}^{d}}$ We must show
+$\tilde{P}_{\mathbb{Q}}(A)=\int_{\mathscr{S}(d)}\tilde{P}_{\omega}(A)d\mathbb{Q}(\omega)\in\{0,1\}$
+Let $P^{*}_{\omega}$ denote the law of the local process $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}$ starting at $\{\omega(x)\}_{x\in\mathbb{Z}^{d}}$.
+Therefore, it is enough to show,
+${P}^{*}_{\mathbb{Q}}(A)=\int_{\mathscr{S}(d)}P^{*}_{\omega}(A)d\mathbb{Q}(\omega)\in\{0,1\}.$
+Denote $\bar{\omega}(n)=\tau_{X_{\Lambda(\omega)}(n)}\omega$.
+Start by noting that under the canonical filtration sequence $\mathscr{G}_{k}=\sigma(\omega,\bar{\omega}(1),\dots,\bar{\omega}(k-1))$, $h(\bar{\omega}(n))$ is a $\tilde{P}_{\mathbb{Q}}$ martingale since
+$\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=\mathbb{E}_{\mathbb{Q}}[\boldsymbol{1}_{A}\circ\theta^{n}|\omega,\bar{\omega}(1),\dots,\bar{\omega}({n})]=h(\bar{\omega}(n))\quad\mathbb{Q}\text{-a.e.}$
 By the Levy Upwards Theorem,
-$h(\bar{\omega}(n))\to\boldsymbol{1}_{A}\qquad(n\to\infty)\quad\mathbb{Q}\text{-a.e.}$ (3)
-Next we show $h=\boldsymbol{1}_{B}$($\mathbb{Q}$ a.e.) for some $B\in\mathcal{B}(d)^{\prime}$. If not then $\mathbb{Q}(h\in[a,b])>0$ for some $0<a<b<1$.
+(2) $h(\bar{\omega}(n))\to\boldsymbol{1}_{A}\qquad(n\to\infty)\quad\mathbb{Q}\text{-a.e.}$
+Next we show $h=\boldsymbol{1}_{B}$($\mathbb{Q}$ a.e.) for some $B\in\mathcal{B}(d)^{\prime}$. If not then
+$\mathbb{Q}(h\in[a,b])>0$ for some $0<a<b<1$.
 But by Birkhoff’s Ergodic Theorem,
@@
 $\mathbb{E}_{\mathbb{Q}}[\Phi]=\mathbb{Q}[h\in[a,b]]>0.$
-But this contradicts 3 since $\Phi=0$ ($\mathbb{Q}$ a.e.). Now we note that from the martingale property and the Markov property and the fact that $\{h(\bar{\omega}(n)\}_{n\geq 0}$ is a martingale,
-$\boldsymbol{1}_{B}(\omega)=\mathbb{E}_{\mathbb{Q}}\Big[\,\boldsymbol{1}_{B}(\bar{\omega}(1))\,\big|\,\omega\,\Big]=K_{d}\,\boldsymbol{1}_{B}(\omega)\quad\mathbb{Q}\text{-a.e.}$ (4)
-Combining 4 with ellipticity,
-$\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{e}\quad(\mu\text{-\text{a.e.} or }\mathbb{Q}\text{-a.e.})\quad\forall\,e\in{V}_{d}.$ (5)
-Iterating 5, we obtain
+But this contradicts 2 since $\Phi=0$ ($\mathbb{Q}$ a.e.). Now we note that from the martingale property and the Markov property, and the fact that $\{h(\bar{\omega}(n))\}_{n\geq 0}$ is a martingale,
+(3) $\boldsymbol{1}_{B}(\omega)=\mathbb{E}_{\mathbb{Q}}\Big[\,\boldsymbol{1}_{B}(\bar{\omega}(1))\,\big|\,\omega\,\Big]=K_{d}^{\Lambda}\,\boldsymbol{1}_{B}(\omega)\quad\mathbb{Q}\text{-a.e.}$
+Combining 3 with ellipticity,
+(4) $\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{e}\quad(\mu\text{-\text{a.e.} or }\mathbb{Q}\text{-a.e.})\quad\forall\,e\in{V}_{d}.$
+Iterating 4, we obtain
 $\boldsymbol{1}_{B}\geq\boldsymbol{1}_{B}\circ\tau_{x}\quad(\mu\text{\text{-a.e.} or }\mathbb{Q}\text{\text{-a.e.}})\quad\forall\,x\in\mathbb{Z}^{d}.$
 Since $\mu[\omega\in B]=\mu[\tau_{x}\omega\in B]$ for every $x\in\mathbb{Z}^{d}$,
-$\boldsymbol{1}_{B}=\boldsymbol{1}_{B}\circ\tau_{x}\quad\mu\text{\text{-a.e.} or}\penalty 10000\ \mathbb{Q}\text{\text{-a.e.}}$ (6)
+(5) $\boldsymbol{1}_{B}=\boldsymbol{1}_{B}\circ\tau_{x}\quad\mu\text{\text{-a.e.} or}\penalty 10000\ \mathbb{Q}\text{\text{-a.e.}}$
 Consider the set
@@
 Observe $(\tau_{x})^{-1}\tilde{B}=\tilde{B}$ which shows $\tilde{B}$ is invariant under shifts.
-From 6 we get that $\mu(\tilde{B})=\mu(B)=0$ or $1$ (by ergodicity).
-From the mutual absolute continuity of $\mathbb{Q}$ and $\mu$, $\mathbb{Q}(B)=\tilde{P}_{\mathbb{Q}}(A)=0$ or $1$.
+From 5 we get that $\mu(\tilde{B})=\mu(B)=0$ or $1$ (by ergodicity).
+From the mutual absolute continuity of $\mathbb{Q}$ and $\mu$, $\mathbb{Q}(B)=P^{*}_{\mathbb{Q}}(A)=0$ or $1$.
 This proves ergodicity of the local process in the environment $\mathbb{Q}$.
+∎
 Thus, in view of the previous proof, we have reduced the proof to the following lemma.
 Lemma 8.5.
-Theorem 3.7 holds under its assumptions if the local process is stationary under some measure $\mathbb{Q}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ which is mutually absolutely continuous to $\mu$.
-We shall develop in the next section the proof of Lemma 8.5.
+Let $\mathbb{Q}$ be a measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic.
+Then $\{\pi_{0}(\tau_{X_{\Lambda(\omega)}(m)}\omega)\}_{m\geq 0}=\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ in distribution (and therefore both are stationary sequences) under $\mathbb{Q}$.
 9. Proof of Lemma 8.5:
-The first part of the proof is to construct $\mathbb{Q}$.
-Let $\lambda^{\prime}$ be an invariant measure of $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 0}$. First, we make an observation on this measure.
-Let $\alpha:\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ given by $\alpha(e_{i})=-e_{i}\quad i=1,2,\dots,d$.
-Lemma 9.1.
-$\lambda^{\prime}$ is an $\alpha$-invariant measure. More precisely,
-$\lambda^{\prime}\circ\alpha(B)=\lambda^{\prime}(B)\quad\forall B\in\mathcal{B}(d)^{\prime}.$
-Proof.
-Start by noting that for every $B\in\mathcal{B}(d)^{\prime}$,
-$\int_{\mathcal{S}(d)}K_{d}^{\Lambda}(\omega,B)d\lambda^{\prime}\circ\alpha=\int_{\mathcal{S}(d)}K_{d}^{\Lambda}(\omega,\alpha(B))d\lambda^{\prime}=\lambda^{\prime}\circ\alpha(B).$
-This shows $\lambda^{\prime}\circ\alpha$ is also an invariant stationary measure for the environment process.
-Furthermore $\lambda^{\prime}\circ\alpha$ is mutually absolutely continuous to $\mu\circ\alpha$.
-One can also show that $\mu\circ\alpha$ is stationary and ergodic with respect to the group-of-shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$.
-Let $g:\mathcal{S}(d)\to\mathbb{R}$ be a measurable function. By Birkhoff’s Ergodic Theorem,
-$\frac{1}{n}\sum_{i=0}^{n-1}g(\tau_{X_{\Lambda(\omega)}(n)}\omega)$ $\to\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\quad\omega\penalty 10000\ \mu\textit{-a.e.}$ (7)
-$\frac{1}{n}\sum_{i=0}^{n-1}g(\tau_{X_{\Lambda(\omega)}(n)}\omega)$ $\to\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\circ\alpha\quad\omega\penalty 10000\ \mu\circ\alpha\textit{-a.e.}$ (8)
-Therefore, we conclude that,
-$\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}=\int_{\mathcal{S}(d)}g(x)d\lambda^{\prime}\circ\alpha,$
-which completes the proof. ∎
-Fix $\omega\in\mathcal{S}(d)$.
+Let $\lambda^{\prime}$ be the stationary ergodic measure of the chain $\{\tau_{X_{\Lambda(\omega)}(n)}\omega\}_{n\geq 0}$. We set $\mathbb{Q}=\lambda^{\prime}$. Fix $\omega\in\mathscr{S}(d)$.
 Recall that $\{X_{\omega}(n)\}_{n\geq 0}$ is the random walk with the law
-$\mathbb{P}[X_{\omega}(1)=x+v\Big|X_{\omega}(0)=x]=\omega(x,v)\quad v\in V_{d}.$ (9)
+(6) $\mathbb{P}[X_{\omega}(1)=x+v\Big|X_{\omega}(0)=x]=\omega(x,v)\quad v\in V_{d}.$
 The environment process is therefore given by $K_{d}^{\operatorname{Id}}\equiv K_{d}$.
 Define the reflected walk, $\{X_{\alpha(\omega)}(n)\}_{n\geq 0}$ with the law
-$\mathbb{P}[X_{\alpha(\omega)}(1)=x+v\Big|X_{\alpha(\omega)}(0)=x]=\omega(x,-v).$ (10)
+(7) $\mathbb{P}[X_{\alpha(\omega)}(1)=x+v\Big|X_{\alpha(\omega)}(0)=x]=\omega(x,-v).$
 The environment process is given by
@@
 $K_{\gamma}(\alpha.\omega,B)=K_{1-\gamma}(\omega,\alpha.B)$.
-This is the control argument using Bernoulli variables as we had mentioned, and now we shall use it explicitly to demonstrate the stationarity of the local process.
-We claim that $\mathbb{Q}=\lambda^{\prime}$.
+This is the control argument using Bernoulli variables as we had mentioned, and now we shall use it explicitly to demonstrate the stationarity of the local process or the equivalent statement of Lemma 8.5.
+Let $\bar{\omega}(k)$ denote the environment process $\tau_{X_{\Lambda(\omega)}(k)}\omega$ for $k\geq 0$.
 Lemma 9.2.
-Let $n\geq 1$.
-Let $B_{0},B_{1},\dots,B_{n}\in\mathcal{B}(d)^{\prime}$.
-Let the probability measure associated with $\omega$ be $\lambda^{\prime}$.
-Let $\bar{\omega}(n)=\tau_{X_{\Lambda(\omega)}(n)}\omega$.
-Then for every $1\leq k\leq n-1$, we have
+Let $n\geq 1$. Let $B_{0},B_{1},\dots,B_{n}\in\mathcal{B}(d)^{\prime}$. Let the probability measure associated with $\omega$ be $\lambda^{\prime}$. Then for every $0\leq k\leq n-1$, we have
 $\lambda^{\prime}\Bigl[\omega\in B_{0},\,\bar{\omega}(1)\in B_{1},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,\gamma_{n}\Bigr]$
-$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\ \bar{\omega}(k+1)\in\alpha(B_{k+1}),\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,1-\gamma_{k+1},\ldots,\gamma_{n}\Bigr].$ (11)
-For $k=0$,
-$\lambda^{\prime}\Big[\omega\in\alpha(B_{0}),\,\bar{\omega}(1)\in B_{1},\,\ldots,\,\bar{\omega}(k)\in B_{k},\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,\gamma_{n}\Bigr]$
-$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\bar{\omega}(1)\in\alpha(B_{1})\ldots,\,\bar{\omega}(k)\in B_{k}\,\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,1-\gamma_{1},\ldots,\gamma_{k},\ldots,\gamma_{n}\Bigr].$ (12)
+$\qquad=\lambda^{\prime}\Bigl[\omega\in B_{0},\,\ldots,\,\bar{\omega}(k)\in\alpha(B_{k}),\ \bar{\omega}(k+1)\in\alpha(B_{k+1}),\ldots,\,\bar{\omega}(n)\in B_{n}\,\Bigm|\,\gamma_{1},\ldots,1-\gamma_{k+1},\ldots,\gamma_{n}\Bigr].$
 To see how this implies stationarity, take $B_{i}=\pi_{0}^{-1}(A_{i})$ where $A_{i}\in\mathcal{B}(d)$. Then
 $\lambda^{\prime}\left[\bigcap_{k=0}^{n}\{\bar{\omega}(k)\in\pi_{0}^{-1}(A_{k})\}\,\middle|\,\gamma_{1},\dots,\gamma_{n}\right]=\frac{1}{2^{n}}\sum_{\{\gamma_{i}\}_{i=1}^{n}\in\{0,1\}^{n}}\lambda^{\prime}\left[\bigcap_{k=0}^{n}\{\bar{\omega}(k)\in\pi_{0}^{-1}(A_{k})\}\,\middle|\,\{\gamma_{i}\}_{1\leq i\leq n}\right]$
-The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{1}),\dots,\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$.
-Now, conclude stationarity using the stationarity of $\bar{\omega}(n)$ and taking $\gamma_{i}=1$.
-It is clear that
+The last expression reduces to $\lambda^{\prime}[\omega\in\pi_{0}^{-1}(A_{0}),\dots,\bar{\omega}(n)\in\pi_{0}^{-1}(A_{n})]$.
+Since $\{\bar{\omega}(k)\}_{k\geq 0}$ is stationary under $\lambda^{\prime}$, this establishes the assertion of Lemma 8.5 and therefore stationarity of $\{\pi_{0}(\tau_{X_{\omega}(k)}\omega)\}_{k\geq 0}$.
+Note that taking $\gamma_{i}=1$, it is clear that
 $\tau_{X_{\Lambda(\omega)}(n)}\omega|\{\gamma_{i}=1\quad i=1,2,\dots,n\}=\tau_{X_{\omega}(n)}\omega.$
+This shows that choosing $\gamma_{i}$’s controls the path suitably.
 Therefore, it is now enough to establish Lemma 9.2.
@@
 $\int_{B_{0}}K_{\gamma}(\alpha.\omega,B_{1})d\lambda^{\prime}\circ\alpha.$
-By Lemma 9.1 and Lemma 9.2, the integral reduces to
-$\int_{B_{0}}K_{1-\gamma}(\omega,\alpha.B_{1})d\lambda^{\prime}\circ\alpha=\lambda^{\prime}(\alpha(\omega)\in B_{0},\bar{\omega}(1)\in\alpha.B_{1}|1-\gamma).$
-Suppose the result is true for $n=m$. Let $n=m+1$. We verify the statement for $k=m+1$ only for simplicity, as others follow from the technique of proof. For $k=m+1$, we write
+By Lemma 9.1, the integral reduces to
+$\int_{B_{0}}K_{1-\gamma}(\omega,\alpha.B_{1})d\lambda^{\prime}\circ\alpha=\mathbb{E}[\boldsymbol{1}_{\omega}(\alpha(B_{0}))\mathbb{E}[\boldsymbol{1}_{\bar{\omega}(1)}(\alpha.B_{1})\Big|\omega,1-\gamma]]$
+This reduces to $\lambda^{\prime}[\omega\in\alpha.B_{0},\bar{\omega}(1)\in\alpha.B_{1}|1-\gamma]$.
+Suppose the result is true for $n=m\geq 1$. Let $n=m+1$. We verify the statement for $k=m+1$ only for simplicity, as others follow from the technique of proof. For $k=m+1$, we write
 $\int_{B_{0}\times B_{1}\times B_{2}\times\dots\times\alpha.B_{m}}K_{\gamma_{m+1}}(\omega_{m},B_{m+1})d\lambda^{\prime}(\omega_{m},\dots,\omega|\gamma_{1},\dots,\gamma_{m}).$
@@
 $\mathbb{E}\Big[\mathbb{E}\Big[\boldsymbol{1}_{\alpha.B_{m+1}}(\bar{\omega}(m+1))\Big|\bar{\omega}(m),1-\gamma_{m+1}\Big]\boldsymbol{1}_{\bar{\omega}(m)}(\alpha.B_{m})\boldsymbol{1}_{\bar{\omega}(m-1)}(B_{m-1})\dots,\boldsymbol{1}_{\omega}(B_{0})\Big|\gamma_{1},\dots,\gamma_{m}\Big].$
-This reduces by definition to the desired form, which completes the proof.
+This reduces by definition to the desired form, which completes the proof of Lemma 9.2.